### v1

# VISTA: Knowledge-Driven Interpretable Vessel Trajectory Imputation via Large Language Models

Abstract.
The Automatic Identification System provides critical information for maritime navigation and safety, yet its trajectories are often incomplete due to signal loss or deliberate tampering. Existing imputation methods emphasize trajectory recovery, paying limited attention to interpretability and failing to provide underlying knowledge that benefits downstream tasks such as anomaly detection and route planning. We propose knowledge-driven interpretable vessel trajectory imputation (VISTA), the first trajectory imputation framework that offers interpretability while simultaneously providing underlying knowledge to support downstream analysis. Specifically, we first define underlying knowledge as a combination of Structured Data-derived Knowledge (SDK) distilled from AIS data and Implicit LLM Knowledge acquired from large-scale Internet corpora. Second, to manage and leverage the SDK effectively at scale, we develop a data–knowledge–data loop that employs a Structured Data-derived Knowledge Graph for SDK extraction and knowledge-driven trajectory imputation. Third, to efficiently process large-scale AIS data, we introduce a workflow management layer that coordinates the end-to-end pipeline, enabling parallel knowledge extraction and trajectory imputation with anomaly handling and redundancy elimination. Experiments on two large AIS datasets show that VISTA is capable of state-of-the-art imputation accuracy and computational efficiency, improving over state-of-the-art baselines by 5%–94% and reducing time cost by 51%–93%, while producing interpretable knowledge cues that benefit downstream tasks. The source code and implementation details of VISTA are publicly available.
PVLDB Reference Format:
PVLDB, 19(1): XXX-XXX, 2026.
doi:$X$.XX/XXX.XX Tianyi Li is the corresponding author.
This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 19, No. 1 ISSN 2150-8097.
doi:$X$.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/hyLiu1994/VISTA.
1. Introduction
The Automatic Identification System (AIS) shares the positions, identities, and other key information of vessels with nearby vessels, coastal stations and satellite, supporting tasks such as route planning, prediction, and anomaly detection (Bank, 2023). However, AIS vessel trajectories are frequently incomplete due to signal loss, equipment failure, or deliberate tampering (Łubkowski and Laskowski, 2017). To address this, trajectory imputation methods have been proposed that employ rule-based interpolation (Widyantara et al., 2023; Zaman et al., 2023; Wang et al., 2023) and advanced deep learning methods (Syed and Ahmed, 2023; Jiang et al., 2024; Zhang et al., 2021). More recently, researchers have recognized the importance of interpretable imputation, aiming to not only recover missing data but also to make the imputation process transparent and understandable to human (Alam et al., 2025; Cen et al., 2024; Fertig et al., 2025).
However, transparency by itself is often insufficient, as downstream tasks may benefit further from access to the underlying knowledge associated with each reconstructed segment: i) an estimation of the behavior pattern supported by concrete evidence, ii) a justification of the imputation result that aligns with the inferred patterns, and iii) a human-friendly explanation that conveys the deeper reasoning underlying the reconstruction of a segment. This idea is illustrated in Figure 1.
Figure 1. Example of knowledge-support imputed trajectory.
Example 0.
AIS loses updates from a cargo ship with “under way using engine” status for about ten minutes near a busy port. The missing sub-trajectory is reconstructed as a smooth curve that connects the observed path before and after the gap. The supporting underlying knowledge can be summarized as follows:
i) Estimation of Behavior Patterns. The behavior corresponds to a decelerate-then-align pattern, frequently observed (about 68%) in AIS data for cargo vessels with queue-following intent near port.
ii) Justification for the Imputation Result. The chosen imputation function (constant turn-rate with smooth speed decay) is most consistent with this pattern, accounting for over 70% of successful reconstructions for the estimated behavior pattern.
iii) Underlying Cause. The behavior reflects port-entry procedures: vessels reduce speed and adjust heading to merge into inbound lanes and approach the pilot boarding line under traffic separation rules.
As illustrated to the lower-left in Figure 1, the underlying cause delivers decision-critical cues (the behavior is explained by port-entry procedures) needed for anomaly detection, sharpening the distinction between expected queue-following behavior and anomalous behavior.
This example highlights a crucial yet underexplored dimension of trajectory imputation: knowledge transfer for downstream tasks. We argue that, in real-world applications, downstream tasks generally will benefit not just from completed trajectories but also the underlying knowledge employed in the imputation process. Such knowledge allows users to enhance downstream tasks by internalizing the behavior patterns, regulatory rules and operational protocols employed during imputation, instead of depending solely on the reconstructed data.
Recent advances in large language models (LLMs) offer promising tools for addressing this challenge. LLMs possess strong capabilities in integrating contextual information, reasoning over domain knowledge, and generating human-understandable explanations (Liang et al., 2025b; Zhang et al., 2024b; Liang et al., 2025a). These properties make them well-suited for knowledge-driven trajectory imputation, where both behavioral plausibility and explanatory value are critical. Yet, applying LLMs in this setting is far from straightforward, as it raises fundamental question about how knowledge should be defined and efficiently extracted, managed, and applied. This leads to three challenges:
C1: How to define underlying knowledge that bridges interpretability and utility? A clear definition of what constitutes underlying knowledge is a prerequisite for its effective extraction, management, and utilization. The knowledge must be general enough to remain human-understandable, and to supply essential cues that can benefit downstream tasks, while also being precise enough to support trajectory imputation. To the best of our knowledge, existing methods (Musleh and Mokbel, 2023; Si et al., 2023; Yu et al., 2025) lack such a systematic formulation of knowledge, often relying on ad hoc heuristics or statistical patterns without explicitly characterizing the knowledge they employ.
C2: How to extract and utilize maritime knowledge effectively at scale? Since LLMs lack maritime domain knowledge (Yang et al., 2025), it is essential to be able to derive such knowledge directly from AIS data. However, AIS records are noisy and heterogeneous, blending meaningful behavioral patterns with irrelevant fluctuations. The key challenge is to extract reliable and interpretable knowledge from raw trajectories and to leverage it effectively to guide imputation and downstream decision-making. This requires mechanisms capable of efficiently organizing, retrieving, and applying maritime knowledge in a structured and scalable manner.
C3: How to efficiently process large-scale AIS data in a well orchestrated manner? While structured management of maritime knowledge is essential, this alone cannot support large-scale deployment. Efficient large-scale AIS processing requires workflows that integrate parallel knowledge extraction and trajectory imputation. Further, LLM inference introduces additional challenges such as stochastic responses, unstable execution times, and redundant or overlapping outputs (Ahn and Yin, 2025; Tan et al., 2025; Agrawal et al., 2024). To address these issues, workflow management must include mechanisms for fault tolerance and de-redundancy to ensure stable and efficient execution across the entire pipeline.
To address these challenges, we propose VISTA, a framework for knowledge-driven interpretable vessel trajectory imputation that provides imputed trajectories together with supporting knowledge.
To address C1, we decompose underlying knowledge into two components: Structured Data-derived Knowledge (SDK) and Implicit LLM Knowledge. SDK, distilled from AIS data, captures vessel attributes, behavior patterns, imputation methods, and their statistical relations, facilitating precision and usability. Implicit LLM Knowledge, learned from large-scale Internet corpora, encodes highly compressed priors. To achieve knowledge that is generalizable for human understanding and yet precise for imputation, SDK serves as a trigger to retrieve and align relevant priors from the LLM, and both are integrated into a unified natural-language format.
To address C2, we establish a data–knowledge–data loop with a Structured Data-derived Knowledge Graph (SD-KG) at its core. The SD-KG serves as a structured and compact representation that organizes vessel attributes, behavior patterns, and imputation methods extracted from AIS data. When going from data to knowledge, VISTA transforms raw AIS data into structured knowledge units by eliminating noise, abstracting behavior patterns, and generating imputation methods, and it then integrates them into the SD-KG. When going from knowledge to data, VISTA leverages the accumulated knowledge in the SD-KG to reconstruct missing trajectories by estimating behaviors and selecting suitable imputation methods guided by evidence. The reconstructed trajectories are then augmented with explanatory context derived from the SD-KG, the goal being to achieve accuracy, interpretability, and applicability to downstream maritime tasks.
To address C3, we design a workflow management layer that enables efficient and reliable processing of large-scale AIS data. It manages the end-to-end pipeline from knowledge extraction to trajectory imputation through coordinated task scheduling, anomaly handling, and redundancy control. This design addresses the inherent instability and inefficiency of LLM inference by introducing mechanisms for parallel execution, fault tolerance, and quality validation, ensuring that large numbers of inference jobs can be executed consistently and efficiently.
The main contributions are summarized as follows:
  • We propose VISTA, a knowledge-driven and interpretable trajectory imputation framework that integrates structured knowledge from AIS data with implicit knowledge from LLMs to enable both reconstruction and explanation.
  • We define underlying knowledge as the combination of Structured Data-derived Knowledge and Implicit LLM Knowledge, and organize it in a Structured Data-derived Knowledge Graph with algorithms for construction, maintenance, and utilization.
  • We design a workflow management layer that coordinates knowledge extraction and trajectory imputation with parallel scheduling, anomaly handling, and redundancy control, enabling efficient and reliable processing of large-scale AIS data.
  • Experiments on two AIS datasets show that VISTA achieves state-of-the-art accuracy, outperforming existing baselines by 5%–94% and reducing inference time by 51%–93%, while producing knowledge cues for downstream applications.
The rest of the paper is structured as follows: Section 2 reviews related work. Section 3 covers notation, the definition of underlying knowledge, and the problem formulation. Section 4 presents VISTA. Section 5 reports experimental results. Section 6 concludes the paper and outlines research directions.
2. Related Work
Table 1 summarizes four kinds of existing trajectory imputation methodologies across three key dimensions: generalization, interpretability, and the ability to provide explicit knowledge. Below we discuss the four major categories in detail.
Rule-based Trajectory Imputation Methods. Prior to the advent of data-driven models, trajectory imputation relied on mathematical rules and physical principles. Classical interpolation methods such as Linear and Cubic Spline Interpolation (Widyantara et al., 2023; Zaman et al., 2023) provide simple, interpretable solutions but ignore geographic or navigational constraints. Extensions like Cubic Hermite Splines incorporate velocity and heading to better capture vessel dynamics, while hybrid methods combine linear and spline interpolation for improved accuracy (Widyantara et al., 2023). Kinematic state estimation methods (e.g., EKF/UKF with CTRA or CMM models (Wang et al., 2023)) enforce consistency with motion laws but remain sensitive to abrupt maneuvers and data sparsity. Overall, rule-based methods are computationally efficient and interpretable, but they lack contextual awareness and generalization to complex maritime environments (Li et al., 2023; Bai et al., 2024; Fu et al., 2017).
Deep Learning-based Trajectory Imputation Methods. With the availability of large-scale AIS data, deep learning has become the dominant paradigm for trajectory imputation. Early works relied on recurrent models such as RNNs, LSTMs, and GRUs, which capture temporal dependencies and overcome the vanishing gradient problem, often combined with clustering strategies to specialize models for distinct shipping routes (Shin and Yang, 2024; Yang et al., 2022; Zhang et al., 2020). Later studies integrated Convolutional Neural Networks with recurrent networks to jointly model local spatial features and temporal evolution (Syed and Ahmed, 2023), while Transformer-based architectures (Wang et al., 2024) now represent the state-of-the-art. Their self-attention mechanism captures long-range dependencies critical for understanding navigational intent. In parallel, Graph Neural Networks (GNNs) (Jiang et al., 2024) have been applied to encode multi-vessel interactions by representing vessels as nodes and their spatial relations as edges, enabling cooperative trajectory prediction in congested waters. Generative methods, particularly GANs (Zhang et al., 2021), Diffusion Models (Zhang et al., 2025; Tashiro et al., 2021; Chen et al., 2024), provide another line of research by learning realistic distributions of vessel motion, producing more diverse and plausible imputations than conventional regression-based methods.
While deep learning methods achieve superior accuracy and strong generalization by directly learning non-linear spatio-temporal patterns, they suffer from poor interpretability. Their decision-making process remains opaque, making it difficult to assess whether outputs are grounded in navigational knowledge or merely reflect statistical correlations. Moreover, these methods do not provide explicit knowledge that can be reused or audited, which limits trust in safety-critical applications (Zhang et al., 2024a).
Table 1. Comparison of existing imputation methods.
Methodology Class Generalization Interpretability Explicit Knowledge
Rule-based Trajectory Imputation No Yes No
Deep Learning-based Trajectory Imputation Yes No No
LLM-based Trajectory Imputation Yes Yes No
Interpretable Trajectory Imputation Yes Yes No
LLM-based Trajectory Imputation Methods. A recent paradigm explores reframing trajectory imputation as a language modeling problem by treating a sequence of GPS points as analogous to words in a sentence. Through spatial discretization into a finite “location vocabulary,” vessel movement can be modeled with a “physical grammar,” enabling the use of Transformer-based architectures such as BERT for trajectory infilling tasks. Representative systems include KAMEL (Musleh and Mokbel, 2023) , which introduces spatially-aware tokenization and multi-point generation to address the limitations of vanilla LLMs, and TrajBERT (Si et al., 2023) , which incorporates spatio-temporal refinement strategies for sparse AIS data. More recent frameworks (e.g., MAKER (Yu et al., 2025) , AIS-LLM (Park et al., 2025) ) extend beyond single-task imputation by fusing trajectories with textual prompts or unifying tasks such as prediction, anomaly detection, and risk assessment, often producing natural language explanations to enhance user trust.
Despite these advances, LLM-based methods still face several fundamental challenges. Tokenizing continuous spatio-temporal data remains non-trivial, generated trajectories are often not physically grounded, and performance can be sensitive to prompt design (Xu et al., 2025). While these models offer stronger process interpretability than deep learning-based methods, they still lack mechanisms to provide explicit, structured knowledge that supports transparent reasoning and consistent integration with downstream tasks.
Interpretable Trajectory Imputation Methods. To overcome the opacity of black-box models in safety-critical maritime systems, interpretable imputation methods aim to balance predictive accuracy with transparency. A prominent direction is the use of Physics-Informed Neural Networks (Alam et al., 2025) , which embed kinematic equations into the loss function so that outputs remain consistent with physical laws. This improves generalization and robustness, particularly under sparse or noisy AIS data. Another line of work (Cen et al., 2024) leverages attention mechanisms to visualize which historical trajectory points influence predictions, providing procedural transparency and aligning model reasoning with domain expertise. Hybrid methods (Fertig et al., 2025) combine neural networks with explicit kinematic models by predicting intermediate physical variables (e.g., acceleration), ensuring that the final trajectory remains physically feasible.
These methods can be seen as either constraining outputs to a “glass box” of physical laws or offering whiteboard-style insights into the model’s internal focus. While such designs enhance both interpretability and generalization compared to pure deep learning, they still lack the ability to provide explicit knowledge that downstream tasks and operators can directly reuse.
3. Preliminaries
3.1. Data and Notation
Definition 0 (AIS Record).
An AIS record $x=(\iota,\lambda,\phi,\tau,\psi,\theta,s,\eta,\chi,\\ d,\ell,\beta,\kappa)$ consists of multiple attributes. At its core, an AIS record is anchored by the spatio-temporal attributes and vessel identifier $(\iota,\lambda,\phi,\tau)$, representing vessel identifier, longitude, latitude, and timestamp, which jointly specify the vessel’s geographic position at a particular time. Other attributes act as auxiliary descriptors:
  • Kinematic: Heading angle $\psi$, course over ground $\theta$, and speed over ground $s$ describe the vessel’s movement.
  • Status-related: Navigation status $\eta$, hazardous cargo type $\chi$, and draught $d$ provide operational and safety context.
  • Static: Vessel length $\ell$, width $\beta$, and type $\kappa$ characterize vessel geometry and category.
Definition 0 (Vessel-specific AIS Record Sequence).
Given a vessel $\iota$, an AIS record sequence $\mathbf{X}_{\iota}=\langle x_{1},x_{2},\dots,x_{T}\rangle$ is a time-ordered sequence, where each $x_{t}$ is as defined in Definition 3.1. The timestamps $\tau$ satisfy $\tau_{t}<\tau_{t+1}$ for all $t\in[1,T-1]$.
A vessel-specific AIS record sequence $\mathbf{X}_{\iota}$ captures the spatio-temporal trajectory of vessel $\iota$, together with its auxiliary descriptors (kinematic, status-related, and static attributes) across the time period $[1,T]$. Given a collection of vessels, the entire AIS data is represented as:
(1) $\mathcal{X}=\{\mathbf{X}_{\iota_{1}},\mathbf{X}_{\iota_{2}},\dots,\mathbf{X}_{\iota_{N}}\},$
where each $\mathbf{X}_{\iota_{i}}$ denotes the AIS record sequence of vessel $\iota_{i}$, and $N$ is the total number of vessels.
3.2. Underlying Knowledge
3.2.1. The definition of underlying knowledge
Definition 0 (Underlying Knowledge in Vessel Trajectory Imputation).
Given AIS data $\mathcal{X}$, the underlying knowledge $\mathcal{U}$ is defined as the integration of:
  • Structured Data-derived Knowledge (SDK) $\mathcal{K}_{d}$, distilled from AIS data $\mathcal{X}$, capturing vessel attributes, behavior patterns, imputation functions, and their empirical dependencies.
  • Implicit LLM Knowledge $\mathcal{K}_{\ell}$, consisting of maritime conventions and tacit operational know-how, together with commonsense knowledge embedded in LLM parameters.
The complete underlying knowledge is then given by
(2) $\mathcal{U}=\Phi(\mathcal{K}_{d},\mathcal{K}_{\ell}),$
where $\Phi$ denotes the process by which SDK $\mathcal{K}_{d}$ serves as trigger to activate and integrate relevant implicit knowledge from $\mathcal{K}_{\ell}$.
As shown in Figure 2, the construction of underlying knowledge $\mathcal{U}$ relies on two complementary sources. First, the SDK $\mathcal{K}_{d}$ is obtained from AIS data $\mathcal{X}$ through LLM-assisted analysis, where entities and their semantic relations can be distilled into a graph structure. Second, the implicit LLM knowledge $\mathcal{K}_{\ell}$ originates from the large-scale Internet corpus. Unlike $\mathcal{K}_{d}$, this knowledge cannot be explicitly enumerated as graph structure, but can be retrieved and aligned through queries triggered by $\mathcal{K}_{d}$. In this way, the structured component $\mathcal{K}_{d}$ serves as a trigger to activate and integrate relevant priors and commonsense information from $\mathcal{K}_{\ell}$, together yielding the complete underlying knowledge $\mathcal{U}$.
Figure 2. Illustration of the underlying knowledge.
3.2.2. The definition of structured data-derived knowledge graph
To enable trajectory imputation to be performed both efficiently and accurately, while also supporting downstream analytics, it is essential to carefully store and manage the SDK $\mathcal{K}_{d}$. Our design is guided by two considerations: 1) Human interpretability and LLM usability, ensuring that the knowledge is not only understandable to domain experts but also directly usable as triggers for LLMs; 2) Imputation accuracy and efficiency, binding vessel- and context-specific information to executable imputation functions with machine-actionable semantics (applicability conditions and parameters), while compressing massive AIS data into compact yet representative knowledge units for scalable retrieval and execution. To address these requirements, we introduce Structured Data-derived Knowledge Graph (SD-KG) to organize SDK.
Definition 0 (Structured Data-derived Knowledge Graph).
The SD-KG is represented as a graph $\mathcal{G}_{d}=(\mathcal{V}_{d},\mathcal{E}_{d})$, distilled from AIS data $\mathcal{X}$, where $\mathcal{V}_{d}$ denotes the set of nodes and $\mathcal{E}_{d}$ denotes their semantic relations. The node set $\mathcal{V}_{d}=\{\mathcal{V}_{s},\mathcal{V}_{b},\mathcal{V}_{f}\}$ is divided into:
  • Static Attribute Nodes $\mathcal{V}_{s}=\{\mathcal{V}_{\iota},\mathcal{V}_{\eta},\mathcal{V}_{\chi},\mathcal{V}_{d},\mathcal{V}_{\ell},\mathcal{V}_{\beta},\mathcal{V}_{\kappa},\mathcal{V}_{\sigma}\}$, mostly derived from the static and status-related attributes of AIS records (Definition 3.1). For each attribute type, a corresponding node set is constructed, where each node represents one possible value of that attribute. Specifically, $\mathcal{V}_{\sigma}$ differs from other sets in $\mathcal{V}_{s}$ as it is derived from core attributes and represents spatial context categories such as shipping lanes, ports, or anchorages.
  • Behavior Pattern Nodes $\mathcal{V}_{b}$ representing characteristic behavior patterns extracted from AIS data. Each node $v_{b}\in\mathcal{V}_{b}$ is defined as a tuple $v_{b}=(p^{s},p^{\theta},p^{\psi},p^{i},p^{\tau})$, where $p^{s}$ denotes the speed pattern, $p^{\theta}$ the course pattern, $p^{\psi}$ the heading pattern, $p^{i}$ the navigation intent, and $p^{\tau}$ the duration time. These patterns are primarily inferred from the kinematic and core attributes.
  • Imputation Function Nodes $\mathcal{V}_{f}$ representing available imputation methods. Each node $v_{f}\in\mathcal{V}_{f}$ is defined as a tuple $v_{f}=(f,d(f))$, where $f$ denotes an imputation function and $d(f)$ denotes its descriptive information. These nodes are primarily linked to the types of behavior patterns they can address, which are inferred from the core attributes of AIS records.
The edge set $\mathcal{E}_{d}=\{\mathcal{E}_{sb},\mathcal{E}_{bf}\}$ consists of two types of relations:
  • Static–Behavior edges $\mathcal{E}_{sb}\subseteq\mathcal{V}_{s}\times\mathcal{V}_{b}$: connecting static attribute nodes and behavior pattern nodes. For an edge $e=(v_{s},v_{b},\\ w)\in\mathcal{E}_{sb}$, the weight $w$ indicates how frequently the static attribute value $v_{s}$ co-occurs with the behavior pattern $v_{b}$ in $\mathcal{X}$.
  • Behavior–Function edges $\mathcal{E}_{bf}\subseteq\mathcal{V}_{b}\times\mathcal{V}_{f}$: connecting behavior pattern nodes and imputation function nodes. For an edge $e=(v_{b},v_{f},w)\in\mathcal{E}_{bf}$, the weight $w$ denotes the frequency with which the imputation method $v_{f}$ successfully imputes trajectories exhibiting the behavior pattern $v_{b}$.
Figure 3. Example of SD-KG.
Example 0.
As shown in Figure 3, consider $\mathcal{G}_{d}=(\{\mathcal{V}_{s},\mathcal{V}_{b},\mathcal{V}_{f}\},\{\mathcal{E}_{sb},\\ \mathcal{E}_{bf}\})$ with static attribute nodes $\mathcal{V}_{s}=\{\mathcal{V}_{\kappa},\mathcal{V}_{\eta}\}$, behavior pattern nodes $\mathcal{V}_{b}=\{v_{b,1},v_{b,2},v_{b,3}\}$, and imputation function nodes $\mathcal{V}_{f}=\{v_{f,1},v_{f,2},v_{f,3}\}$. The pattern $v_{b,3}=(p^{s},p^{\theta},p^{\psi},p^{i},p^{\tau})$ corresponds to stable speed ($p^{s}$), gradual course change ($p^{\theta}$), low heading fluctuation ($p^{\psi}$), lane-following intent ($p^{i}$), and duration 1300–1350s ($p^{\tau}$). The edge $(v_{\eta,3},v_{b,3},6)\in\mathcal{E}_{sb}$ (red edge in Figure 3) shows that vessels with status “under way using engine” co-occurred with this pattern six times, while $(v_{b,3},v_{f,1},6)\in\mathcal{E}_{bf}$ (blue edge in Figure 3) indicates that the Linear Interpolation function $v_{f,1}=(f,d(f))$ successfully imputed this pattern six times.
This design guarantees the key requirements identified earlier.
Human interpretability and LLM usability. The construction achieves interpretability by explicitly organizing knowledge into semantically meaningful node types—static attributes, behavioral patterns, and imputation functions, that are readily understood by domain experts, while also providing deterministic structures that serve as precise triggers for $\mathcal{K}_{l}$. The decomposition of behaviors into speed, course, heading, intent, and duration enhances transparency, and the explicit pairing of imputation functions with descriptive semantics enables both humans and LLMs to reason about why a given function is appropriate for a specific pattern.
Imputation accuracy and efficiency. Accuracy is ensured in two complementary ways. First, the set of imputation functions $\mathcal{V}_{f}$ is derived directly from real AIS data, capturing empirically validated and executable strategies that can be invoked to reproduce vessel behaviors under similar conditions. Second, edge weights quantify the statistical association between attributes, patterns, and functions, thereby guiding the selection of imputation methods with data-grounded priors rather than unconstrained heuristics. Efficiency is achieved by compressing massive AIS data into a compact, executable knowledge graph representation, enabling inference through lightweight subgraph retrieval and direct function execution, while supporting incremental updates without retraining.
Together, these mechanisms ensure that $\mathcal{G}_{d}$ provides an accurate, efficient, and interpretable foundation for knowledge-driven vessel trajectory imputation. Further details on the construction and maintenance of SD-KG are provided in Section 4.2.
Figure 4. Partition of the AIS record sequence $\mathbf{X}_{\iota}$ into minimal segments (4 records per segment).
3.3. Problem Definition
We formalize the problem in three parts. With AIS records defined (Definitions 3.13.2), we first capture data incompleteness via minimal segments and an observation mask. Next, for complete AIS data, each segment is distilled into a knowledge unit and used to update the SD-KG $\mathcal{G}_{d}$. Finally, for incomplete segments, We define knowledge-driven trajectory imputation that queries the SD-KG $\mathcal{G}_{d}$ under contextual constraints to retrieve candidate behavior patterns and imputation functions; and then we selects and applies the most plausible function to reconstruct the segment, and returns the imputed segment together with justification composed of SD-KG statistics and grounded rationales.
3.3.1. Incomplete AIS Data
Trajectory missing in AIS data typically occurs in the form of block missing (Łubkowski and Laskowski, 2017), where consecutive records are absent rather than isolated points. To capture this phenomenon in a structured way, we introduce the concept of a minimal segment, which serves as the atomic block for modeling both observed and missing intervals.
Definition 0 (Minimal Segment).
For a vessel-specific AIS record sequence $\mathbf{X}_{\iota}=\langle x_{1},x_{2},\dots,x_{T}\rangle$, a minimal segment $\mathbf{S}_{[t,t+m-1]}$ is defined as a consecutive block of AIS records
(3) $\mathbf{S}_{[t,t+m-1]}=\langle x_{t},x_{t+1},\dots,x_{t+m-1}\rangle,$
where $m$ denotes the fixed segment length, corresponding to the minimal unit for modeling missing trajectory data (default $m=20$).
Thus, an AIS record sequence $\mathbf{X}_{\iota}=\langle x_{1},x_{2},\dots,x_{T}\rangle$ can be uniquely partitioned into an ordered concatenation of disjoint minimal segments
(4) $\mathbf{X}_{\iota}=\mathbf{S}_{\iota}^{1}\,\|\,\mathbf{S}_{\iota}^{2}\,\|\,\dots\,\|\,\mathbf{S}_{\iota}^{K},$
where each $\mathbf{S}_{\iota}^{k}=\mathbf{S}_{[t_{k},\,t_{k}+m-1]}$ denotes the $k$-th minimal segment of fixed length $m$, ensuring that all segments are non-overlapping and together cover the entire AIS sequence of vessel $\iota$.
Figure 5. Example of updating the SD-KG.
Example 0.
As shown in Figure 4, let the fixed segment length be $m=4$. Around index $t$, the AIS record sequence $\mathbf{X}_{\iota}$ is partitioned into three consecutive minimal segments: $\mathbf{S}_{\iota}^{k-1}=\langle x_{t-4},x_{t-3},x_{t-2},x_{t-1}\rangle$, $\mathbf{S}_{\iota}^{k}=\langle x_{t},x_{t+1},x_{t+2},x_{t+3}\rangle$, and $\mathbf{S}_{\iota}^{k+1}=\langle x_{t+4},x_{t+5},x_{t+6},x_{t+7}\rangle$.
Definition 0 (Observation Mask).
Given the fixed-length partition $\mathbf{X}_{\iota}=\mathbf{S}_{\iota}^{1}\,\|\,\mathbf{S}_{\iota}^{2}\,\|\,\dots\,\|\,\mathbf{S}_{\iota}^{K}$, the segment mask is the binary vector
(5) $\mathbf{M}_{\iota}=\langle M_{1},\dots,M_{K}\rangle,$
where
[…]
Example 0.
Continuing Example 3.7, using the convention $M_{k}=0$ for a segment with missing records and $M_{k}=1$ for a fully observed segment, the configuration in Figure 4 yields $M_{k-1}=1$, $M_{k}=0$, and $M_{k+1}=0$, i.e., the local portion of the mask vector $\mathbf{M}_{\iota}$ is $\langle\ldots,1,0,0,\ldots\rangle$.
3.3.2. Updating of Structured Data-derived Knowledge
A minimal segment $\mathbf{S}_{\iota}^{k}$ is a time-indexed sample and cannot directly update the SD-KG $\mathcal{G}_{d}$, which is organized over semantic node types and their relations. To bridge segment-level data and the graph-level representation, we introduce the knowledge unit as the minimal semantic unit distilled from a segment: it aligns static attribute, behavior pattern, and imputation function, serving as the atomic granularity for SD-KG updating.
Definition 0 (Knowledge Unit).
A knowledge unit $u$ is defined as a triple $u\;=(v_{s},v_{b},v_{f})\;=((\iota,\eta,\chi,d,\ell,\beta,\kappa,\sigma),\;(p^{s},p^{\theta},p^{\psi},p^{i},p^{\tau}),\;\\ (f,d(f)))$, where $v_{s}$ is static attributes, $v_{b}$ is a behavior pattern, and $v_{f}$ is an imputation method.
On this basis, we formally define how $\mathcal{G}_{d}$ is incrementally updated from complete segments.
Definition 0 (Updating of Structured Data-derived Knowledge Graph).
Given a minimal segment $\mathbf{S}_{\iota}^{k}$ with $M_{\iota}^{k}=1$, the updating of SD-KG refers to the procedure
(7) $\mathcal{G}_{d}\;\leftarrow\;\mathrm{Update}(\mathcal{G}_{d},u_{\iota}^{k}),$
where $u_{\iota}^{k}$ is the knowledge unit distilled from $\mathbf{S}_{\iota}^{k}$. Formally, the update consists of two steps:
  • Knowledge distillation: $\mathbf{S}_{\iota}^{k}\mapsto u_{\iota}^{k}=(v_{s},v_{b},v_{f})$, which compresses the raw segment into static attributes $v_{s}$, behavior pattern $v_{b}$, and imputation method $v_{f}$.
  • Graph integration and maintenance: The node sets $\mathcal{V}_{d}$ and edge sets $\mathcal{E}_{d}$ of $\mathcal{G}_{d}$ are updated by inserting or merging $v_{s},v_{b},v_{f}$ and adjusting the corresponding edges $(v_{s},v_{b})\in\mathcal{E}_{sb}$ and $(v_{b},v_{f})\in\mathcal{E}_{bf}$. Redundant nodes are merged and edge weights aggregated to ensure that $\mathcal{G}_{d}$ remains compact and efficient for retrieval.
Example 0.
As illustrated in Figure 5, there is a segment $\mathbf{S}_{\iota}^{k}=\langle x_{t},\ldots,x_{t+m-1}\rangle$ with $M_{\iota}^{k}=1$. Within this segment, the speed $s$ remains stable at approximately $5.4$ kn, the course $\theta$ varies gradually around $110^{\circ}$, the heading $\psi$ exhibits mild fluctuations near $110^{\circ}$, and both the navigation status and vessel type are constant, recorded as “under way using engine” and “cargo”.
Step 1: Knowledge distillation. From the static and status-related attributes, together with the vessel’s geographic position, we derive the static attributes $v_{s}=(\iota,\eta,\kappa,\sigma)$, describing a cargo vessel $\iota$ operating “under way using engine” near the Traffic Separation Scheme in the Great Belt East Channel. From the kinematic attributes, we extract the behavior pattern $v_{b}=(p^{s},p^{\theta},p^{\psi},p^{i},p^{\tau})$, corresponding to stable speed, gradual course change, mild heading fluctuation, lane-following intent, and duration 1300130013501350s. To avoid redundancy in $\mathcal{V}_{b}$, the states $p^{s},p^{\theta},p^{\psi},p^{i},p^{\tau}$ are preferentially aligned with existing entries, in this case matching $v_{b,3}\in\mathcal{V}_{b}$ (continuing Example 3.5). Finally, based on the trajectory trend, we generate the imputation method $v_{f}=(f,d(f))$, where $f$ is the execute function that simulates the observed dynamics and $d(f)$ records its description.
Step 2: Graph integration and maintenance. Continuing Example 3.5, $v_{b}$ coincides with $v_{b,3}$ and the imputation function $v_{f}$ is semantically equivalent to $v_{f,1}$. The graph is updated accordingly: the weights of edges $(v_{\eta,3},v_{b,3})$ and $(v_{\kappa,2},v_{b,3})$ are each incremented by $1$ (from $6$ to $7$, red edges), while a new edge $(v_{b,3},v_{f,1})$ is added with weight $1$ (blue edge).
Further details on the updating mechanism of SD-KG are presented in Section 4.2.
3.3.3. Knowledge-Driven Interpretable Trajectory Imputation
Before specifying the imputation procedure, we first formalize the target output. The outcome should not only reconstruct the missing trajectory but also incorporate supporting knowledge that justifies the decision process. This leads to the following definition.
Definition 0 (Knowledge-supported Imputation Outcome).
For a minimal segment $\mathbf{S}_{\iota}^{k}$ with $M_{\iota}^{k}=0$, a knowledge-supported imputation outcome is defined as $R_{\iota}^{k}=(\widehat{\mathbf{S}}_{\iota}^{k},\;\mathcal{J}_{\iota}^{k})$ where $\widehat{\mathbf{S}}_{\iota}^{k}$ denotes the imputed trajectory segment, and $\mathcal{J}_{\iota}^{k}$ represents the accompanying justification. The justification $\mathcal{J}_{\iota}^{k}$ consists of:
  • Behavior-pattern rationale $(v_{b}^{*},\,\mathcal{J}_{\iota}^{k,b})$: $v_{b}^{*}$ is the selected behavior pattern for $\mathbf{S}_{\iota}^{k}$, and $\mathcal{J}_{\iota}^{k,b}$ provides contextual and statistical evidence from $\mathcal{G}_{d}$ supporting this selection.
  • Function-selection rationale $(v_{f}^{*},\,\mathcal{J}_{\iota}^{k,f})$: $v_{f}^{*}$ is the selected imputation function applied to obtain $\widehat{\mathbf{S}}_{\iota}^{k}$, and $\mathcal{J}_{\iota}^{k,f}$ contains statistical evidence from $\mathcal{G}_{d}$ supporting the choice of this function.
  • Human-friendly explanation $\mathcal{J}_{\iota}^{k,h}$: a natural-language interpretation, which provides high-level reasoning on why the vessel may exhibit the chosen behavior.
Example 0.
Continuing Example 1.1, the imputed segment is $\widehat{\mathbf{S}}_{\iota}^{k}$, reconstructed as a smooth curve bridging the missing interval, as shown in upper-left in Figure 1. The selected behavior pattern $v_{b}^{*}$ is a decelerate-then-align maneuver with low heading fluctuation and queue-following intent, while its supporting evidence $\mathcal{J}_{\iota}^{k,b}$ indicates that this pattern occurs in about $68\%$ of comparable gaps in historical AIS data. The chosen imputation function $v_{f}^{*}$ is “CTR+SD”, and $\mathcal{J}_{\iota}^{k,f}$ records statistical support that over $70\%$ of such gaps are best reconstructed by this function. Finally, $\mathcal{J}_{\iota}^{k,h}$ provides a natural-language explanation: the maneuver arises from port-entry procedures, where vessels reduce speed and adjust heading to merge into inbound lanes under traffic separation rules.
Definition 0 (Knowledge-Driven Trajectory Imputation).
Given a minimal segment $\mathbf{S}_{\iota}^{k}$ with $M_{\iota}^{k}=0$ and contextual knowledge units $(u_{\iota}^{-},u_{\iota}^{+})$, the knowledge-driven trajectory imputation produces a knowledge-supported imputation outcome
(8) $R_{\iota}^{k}\;\leftarrow\;\mathrm{Impute}(\mathbf{S}_{\iota}^{k},(u_{\iota}^{-},u_{\iota}^{+}),\mathcal{G}_{d},\mathcal{K}_{\ell}),$
where $u_{\iota}^{-}$ and $u_{\iota}^{+}$ denote the contextual knowledge units extracted from the nearest complete segments preceding and succeeding $\mathbf{S}_{\iota}^{k}$, respectively. The procedure consists of:
  • Behavior pattern selection and rationale: Using the contextual knowledge units $(u_{\iota}^{-},u_{\iota}^{+})$ as query constraints, $\mathcal{G}_{d}$ is searched to retrieve a candidate set $\mathcal{C}_{b}\subseteq\mathcal{V}_{b}$. A reasoning process then selects the most plausible behavior pattern $v_{b}^{*}\in\mathcal{C}_{b}$, and the retrieved set with associated evidence is recorded as $\mathcal{J}_{\iota}^{k,b}$.
  • Imputation method selection, execution and rationale: Using the selected behavior pattern $v_{b}^{*}$ as query constraints, $\mathcal{G}_{d}$ is queried to retrieve $\mathcal{C}_{f}\subseteq\mathcal{V}_{f}$. An imputation procedure selects the most suitable method $v_{f}^{*}\in\mathcal{C}_{f}$ and applies it to obtain $\widehat{\mathbf{S}}_{\iota}^{k}$. The retrieved set and supporting statistics are recorded as $\mathcal{J}_{\iota}^{k,f}$.
  • Human-friendly explanation: Given the selected behavior pattern and imputation function $(v_{b}^{*},v_{f}^{*})$, along with the induced subgraph $\mathcal{G}_{d}[\mathcal{V}_{s}(\iota)\cup\{v_{b}^{*}\}\cup\{v_{f}^{*}\}]$ of $\mathcal{G}_{d}$ (where $\mathcal{V}_{s}(\iota)$ denotes vessel-specific statistics), an explanatory generation step produces the natural-language justification $\mathcal{J}_{\iota}^{k,h}$, providing high-level reasoning for the reconstructed behavior.
Further details on knowledge driven trajectory imputation are provided in Section 4.3.
Figure 6. Overview of VISTA.
4. Knowledge-Driven Interpretable Vessel Trajectory Imputation via LLMs
4.1. Overview of VISTA
As illustrated in Figure 6, VISTA establishes a closed-loop that transforms raw AIS data into structured maritime knowledge and reuses this knowledge to reconstruct missing trajectories with interpretable reasoning. To ensure scalable and reliable execution across large-scale AIS datasets, VISTA is supported by a workflow management layer that enables parallel processing, anomaly handling, and redundancy control.
Data-Knowledge-Data Loop. The core loop of VISTA consists of two interconnected stages: SD-KG Construction and Maintenance and Knowledge-Driven Trajectory Imputation. The first stage converts raw AIS data into structured knowledge through static and spatial encoding, behavior abstraction, and imputation method generation. Each resulting knowledge unit, consisting of vessel attributes, behavior patterns, and validated imputation functions, is integrated into the SD-KG via node alignment and edge-weight updates, gradually building a compact and updatable knowledge repository. The second stage reuses this knowledge to reconstruct incomplete trajectories. It estimates behavior patterns with graph-supported evidence, selects and executes suitable imputation functions, and composes explanations that link reconstructed behaviors to maritime rules and operational rationales.
Workflow Manager Layer. To handle the computational demands of large-scale AIS data and LLM-based inference, VISTA incorporates a dedicated workflow management layer operating in parallel with both stages. It includes two specialized managers: the SD-KG Construction Workflow Manager, which orchestrates parallel knowledge extraction, anomaly handling, and de-redundancy, and the Trajectory Imputation Workflow Manager, which manages concurrent imputation and runtime anomaly guard.
4.2. SD-KG Construction and Maintenance
We construct and maintain the SD-KG through three modules followed by a graph update. For each complete minimal segment, the Static and Spatial Encoder produces static attributes and spatial context, the Behavior Abstraction maps motion signals to a behavior pattern, and the Imputation Method Builder registers or generates a validated executable method with a concise description. The outputs form a knowledge unit $u=\langle v_{s},v_{b},v_{f}\rangle$ as defined in Definition 3.10. Each knowledge unit is then integrated into the SD-KG through node alignment, and edge updates.
4.2.1. Static and Spatial Encoder
The encoder organizes the static attribute $v_{s}=\langle\iota,\eta,\chi,d,\ell,\beta,\sigma,\kappa\rangle$ into three classes and applies targeted processing. Discrete fields $\{\iota,\eta,\chi,\kappa\}$ are standardized, continuous fields $\{d,\ell,\beta\}$ are discretized into intervals, and spatial context $\{\sigma\}$ is inferred from geospatial surroundings. This design keeps the $\mathcal{V}_{s}$ in SD-KG compact, interpretable, and easy to maintain.
Discrete Attributes. Vessel identifier $\iota$, navigation status $\eta$, hazardous cargo type $\chi$, and vessel type $\kappa$ are directly obtained from the corresponding static attributes of AIS records $x$ within the minimal segment $\mathbf{S}_{\iota}^{k}$ and standardized to canonical vocabularies.
Continuous Attributes. Draught $d$, length $\ell$, and width $\beta$ are discretized into interval-based categories to form finite vocabularies $\mathcal{V}_{d},\mathcal{V}_{\ell},\mathcal{V}_{\beta}$ and preserve interpretability. Specifically, $d$ is discretized into 2 m intervals up to 12 m, $\ell$ into 50 m intervals up to 300 m, and $\beta$ into 5 m intervals up to 30 m, each with an open-ended bin beyond the upper bound. These cutoffs reflect common maritime conventions (Authority, 2022; Clarksons, 2025) (e.g., Panamax thresholds at $\ell\approx 300$ m, $\beta\approx 32$ m, and $d\approx 12$ m) while maintaining sufficient support.
Spatial Context. The surrounding maritime environment is inferred by querying external sources based on coordinates $(\lambda,\phi)$ in AIS records within the minimal segment $\mathbf{S}_{\iota}^{k}$ to detect nearby infrastructures such as traffic separation schemes, shipping lanes, or anchorages (e.g., Overpass API (Overpass API Contributors, 2025)). The most frequent category within $\mathbf{S}_{\iota}^{k}$ is selected as the representative $\sigma$.
4.2.2. Behavior Abstraction
This component converts motion signals within the minimal segment into a discrete, human interpretable behavior pattern $v_{b}=(p^{s},p^{\theta},p^{\psi},p^{i},p^{\tau})$ and constrains vocabulary growth for interpretability and efficient retrieval.
Kinematic Tokens $(p^{s},p^{\theta},p^{\psi})$. Speed $s$, course over ground $\theta$, and heading $\psi$ exhibit diverse dynamics; direct tokenization yields opaque categories and uncontrolled growth. An LLM assisted procedure analyzes the temporal evolution in $\mathbf{S}_{\iota}^{k}$ and maps them to discrete variables $(p^{s},p^{\theta},p^{\psi})$. Vocabulary size is controlled by finite sets $\mathcal{P}^{s},\mathcal{P}^{\theta},\mathcal{P}^{\psi}$ with a deduplication rule that reuses semantically equivalent tokens before adding new ones.
Intent Token ($p^{i}$). Navigation intent $p^{i}$ is inferred from static attributes $v_{s}$ and contextual cues $(p^{s},p^{\theta},p^{\psi},p^{\tau})$ through LLM analysis. We maintain a finite intent set $\mathcal{P}^{i}$ with the same deduplication rule; ambiguous cases fall back to an undetermined intent.
Duration Token ($p^{\tau}$). Duration token $p^{\tau}$ is computed from the segment time span and discretized into fixed width 50 s intervals, with an open ended bin for long gaps (i.e., $3600,+\infty)$).
4.2.3. Imputation Method Builder
The Imputation Method Builder produces the imputation method $v_{f}=(f,d(f))$, where $f$ is executable function and $d(f)$ is a concise usage description. Given a minimal segment $\mathbf{S}_{\iota}^{k}$, $v_{f}$ is generated through a three-stage process.
Function Proposal. Based on $\mathbf{S}_{\iota}^{k}$ and the associated $(v_{s},v_{b})$, we first retrieve the most relevant imputation method $v_{f}\in\mathcal{V}_{f}$ using the retrieval scheme described in Section 4.3.2, to prevent unnecessary duplication and keep $\mathcal{V}_{f}$ compact. If no suitable candidate is found, a LLM-based function generator produces a new executable function $f$ that can be directly applied to fit the observed trajectory within the segment.
Execution and Validation. The proposed function $f$ is executed on $\mathbf{S}_{\iota}^{k}$. At each timestamp we compute the absolute errors in latitude and longitude, then average them to obtain the overall fitting error $e(f)=\tfrac{1}{2}(\mathrm{MAE}_{\phi}+\mathrm{MAE}_{\lambda})$. The function is accepted if $e(f)$ does not exceed a predefined threshold of $\varrho_{f}=3e-3$. If the error exceeds this threshold, LLM-based function generator is provided with diagnostic feedback (error summaries), and asked to refine its proposal. This is repeated at most $\varrho_{r}$ times (default 3).
Description and Alignment. For each accepted function, if no descriptive element $d(f)$ is available, the LLM-based function explainer generates $d(f)$ conditioned on the executable code $f$ and the selected behavior pattern $v_{b}$. The description specifies the intended use, key assumptions, and parameter definitions, thereby improving the interpretability of the imputation method.
4.2.4. SD-KG Integration and Maintenance
Once knowledge units are extracted, the SD-KG is incrementally updated through two operations: node integration and edge maintenance.
Node Integration. Since static attributes, behavior patterns, and imputation methods have already undergone de-redundancy during knowledge unit extraction, the integration process is straightforward. For each static attribute, if a newly generated token does not yet exist, a corresponding node is created and added to $\mathcal{V}_{s}$. Similarly, if the behavior pattern $v_{b}$ and $v_{f}$ is not present in $\mathcal{V}_{b}$ and $\mathcal{V}_{f}$, a new node is created and inserted.
Edge Maintenance. For each knowledge unit $u=\langle v_{s},v_{b},v_{f}\rangle$, edges $(v_{s},v_{b})$ and $(v_{b},v_{f})$ are updated. If the edge does not exist, it is created with an initial weight of $1$. If it already exists, the weight is incremented by $1$, capturing the frequency of observed co-occurrences. This ensures that the SD-KG preserves both structural completeness and statistically grounded associations while remaining compact and retrievable.
Details of the LLM prompts and the overall SD-KG construction and maintenance procedure are provided in Appendix A.1.
4.3. Knowledge-Driven Trajectory Imputation
With the SD-KG $\mathcal{G}_{d}$ in place, we formalize knowledge driven trajectory imputation for an incomplete minimal segment $\mathbf{S}_{\iota}^{k}$. The procedure comprises three components: Behavior Estimator, which uses contextual constraints to query $\mathcal{G}_{d}$, rank candidates, and select the behavior pattern with supporting evidence; Method Selector, which retrieves an imputation method aligned with the chosen pattern, and executes it to reconstruct the gap; and Explanation Composer, which derives a human friendly explanation by combining a regulatory rule cue with an operational protocol rationale from the induced subgraph of SD-KG and contextual informations. In what follows, we detail each component.
4.3.1. Behavior Estimator
It operates in two steps: Candidate Set Generation, which queries the SD-KG to produce a ranked shortlist of plausible behaviors, and LLM-driven Selection and Rationale, which chooses the final pattern and explains the choice.
Candidate Set Generation. Given an incomplete segment $\mathbf{S}_{\iota}^{k}$ with context $(u_{\iota}^{-},u_{\iota}^{+})$, it infers the vessel-specific static attribute set $\mathcal{V}^{k}_{s}(\iota)$ from $(u_{\iota}^{-},u_{\iota}^{+})$, due to static attributes vary slowly over time. It then queries the SD-KG to retrieve behavior candidates via static–behavior edges
(9) $\mathcal{C}_{b}\;=\;\{\,v_{b}\in\mathcal{V}_{b}\ |\ \sum_{v\in\mathcal{V}^{k}_{s}(\iota)}w_{sb}(v,v_{b})>0\,\},$
and assign each candidate a graph prior based on normalized multiplicative support
(10) $\pi(v_{b})\;=\;\frac{\displaystyle\prod_{v\in\mathcal{V}^{k}_{s}(\iota)}\!(w_{sb}(v,v_{b})+1)}{\displaystyle\sum_{v_{b}^{\prime}\in\mathcal{C}_{b}}\ \prod_{v\in\mathcal{V}^{k}_{s}(\iota)}\!(w_{sb}(v,v_{b}^{\prime})+1)},$
where $w_{sb}$ is the weight in edge $(v,v_{b},w)$. It keeps the top-$K$ candidates by $\pi(v_{b})$ as the shortlist $\mathcal{C}_{b}^{K}$ (default K=5K{=}5).
LLM-driven Selection and Rationale. It first constructs the induced subgraph $\mathcal{G}_{\iota,b}^{K}=\mathcal{G}_{d}[\mathcal{V}^{k}_{s}(\iota)\cup\mathcal{C}_{b}^{K}]$, serializes it in a graph description language (i.e., DOT (Wikipedia contributors, 2025a)), and supplies it together with the contextual behavior patterns $v_{b}^{-}$ and $v_{b}^{+}$ extracted from $u_{\iota}^{-}$ and $u_{\iota}^{+}$. It then selects a final pattern $v_{b}^{*}\in\mathcal{C}_{b}^{K}$ based on the LLM’s analysis of these information.
Alongside the selection, it also obtains a two-part rationale: i) graph support, which highlights the most informative edges in $\mathcal{G}_{\iota,b}^{K}$ and their weights that favor the choice; and ii) contextual justification, which explains the consistency with $v_{b}^{-}$, $v_{b}^{+}$, and the gap’s boundary conditions. These two items constitute $\mathcal{J}_{\iota}^{k,b}$.
4.3.2. Method Selector
This component operates similarly to the Behavior Estimator but omits contextual information during selection, as the selected behavior $v_{b}^{*}$ already encapsulates the relevant kinematic and intent information. Given the selected behavior $v_{b}^{*}$, it retrieves candidate functions through the behavior–function edges and scores them using the same prior construction as in Equations 9 and 10, yielding a ranked shortlist $\mathcal{C}_{f}^{K}$. It then builds the induced subgraph $\mathcal{G}_{\iota,f}^{K}=\mathcal{G}_{d}[\{v_{b}^{*}\}\cup\mathcal{C}_{f}^{K}]$, serialize it in DOT, and select the best method $v_{f}^{*}\in\mathcal{C}_{f}^{K}$ by leveraging the LLM’s analysis of graph evidence and compatibility with the chosen behavior $v_{b}^{*}$. Next, it executes $f\in v_{b}^{*}$ to obtain the reconstruction $\widehat{\mathbf{S}}_{\iota}^{k}$. Alongside the selection, it records graph support by citing the informative edges within $\mathcal{G}_{\iota,f}^{K}$, which forms the rationale $\mathcal{J}_{\iota}^{k,f}$.
4.3.3. Explanation Composer
Explanation composer delivers a human-friendly explanation designed for downstream tasks to internalize two kinds of domain knowledge: a Regulatory-rule cue (under which rules), and an Operational-protocol rationale (why it typically happens here). It analyzes a compact evidence view composed of the selected behavior $v_{b}^{*}$, the selected method $v_{f}^{*}$, static attributes $\mathcal{V}_{s}(\iota)$, spatial context $\sigma$, boundary patterns $(v_{b}^{-},v_{b}^{+})$, and the induced subgraph $\mathcal{G}_{d}[\mathcal{V}_{s}(\iota)\cup\{v_{b}^{*}\}\cup\{v_{f}^{*}\}]$. Specifically, the explanation consists of the following components:
Regulatory-rule Cue. A minimal yet sufficient indication of the governing navigation or traffic rule in context, expressed as a rule label with applicability conditions and a spatial anchor (e.g., “TSS inbound-lane compliance; applies at port-entrance traffic separation; near pilot boarding line”). It is activated by the spatial context $\sigma$ (e.g., TSS, anchorage, pilot line, speed-restricted area), static vessel attributes (e.g., vessel type, vessel size), and the intent in $v_{b}^{*}$. The cue is generated in a concise, human-readable form, highlighting only the most relevant rules for the current scenario.
Operational-protocol Rationale. A procedural-level explanation of why the selected behavior pattern typically occurs in this context and how it relates to the corresponding rule cue. It emphasizes operational logic such as sequencing, merging, pilot approach, collision avoidance, or weather avoidance. By semantically aligning $v_{b}^{*}$ with the rule cue and spatial context $\sigma$, the reasoning process infers the most plausible operational routine and briefly contrasts it with salient alternatives (e.g., explaining why this is not a swerve for collision avoidance). The rationale remains independent of the imputation method, focusing instead on institutional and procedural logic that underpins vessel operations.
Further details (i.e., prompts and overall procedure) are provided in Appendix A.2.
Table 2. Overall effectiveness comparison on AIS-DK and AIS-US datasets.
Method AIS-DK AIS-US
MAE(Lat) RMSE(Lat) MAE(Lon) RMSE(Lon) MHD MAE(Lat) RMSE(Lat) MAE(Lon) RMSE(Lon) MHD
Lin-ITP 2.650e-3 6.386e-2 2.708e-3 5.379e-2 0.3978 1.497e-2 1.453e-1 3.336e-2 4.277e-1 3.9662
Akima Spline 2.127e-3 5.984e-2 2.188e-3 4.833e-2 0.3195 1.274e-2 1.510e-1 3.070e-2 4.539e-1 3.5812
Kalman Filter 2.026e-3 5.664e-2 2.158e-3 4.799e-2 0.3085 1.250e-2 1.294e-1 2.742e-2 3.848e-1 3.2864
Multi-task AIS 1.992e-3 5.364e-2 2.086e-3 4.692e-2 0.2932 1.113e-2 1.164e-1 2.593e-2 3.684e-1 3.1316
MH-GIN 1.906e-3 4.923e-2 1.895e-3 4.417e-2 0.2836 8.728e-3 8.765e-2 2.231e-2 3.292e-1 2.2164
KAMEL 1.975e-3 5.123e-2 1.950e-3 4.539e-2 0.2875 9.875e-3 9.123e-2 2.450e-2 3.439e-1 2.6268
Qwen-plus-th 5.208e-3 5.734e-2 8.694e-3 5.063e-2 0.9139 8.132e-2 2.599e+0 3.027e-2 3.751e-1 10.9758
Qwen-flash-th 2.161e-2 5.622e-1 1.402e-1 3.852e+0 9.5032 7.106e-2 8.314e-1 1478e+1 6.091e+1 482.6334
GLM-4.5-th 2.641e-3 4.929e-2 3.675e-3 1.095e-2 0.3278 9.896e-3 9.431e-2 2.464e-2 3.467e-1 2.9531
GLM-4.5-air-th 6.265e-3 6.201e-2 8.738e-3 5.122e-2 1.0216 2.301e-2 1.579e-1 3.075e-2 3.970e-1 4.9803
VISTA 1.817e-3 4.324e-2 1.123e-3 4.027e-2 0.2418 4.388e-3 1.455e-2 5.691e-3 2.052e-2 0.7945
Table 3. Overall time cost comparison.
Dataset Qwen-plus-th Qwen-flash-th GLM-4.5-th GLM-4.5-air-th VISTA
AIS-DK 30:15:11 14:14:04 91:07:30 25:07:30 6:32:37
AIS-US 24:09:11 12:10:08 88:52:08 22:22:52 6:01:11
4.4. Workflow Management Layer
While the previous sections establish how VISTA extracts knowledge and performs trajectory imputation, both stages rely heavily on LLM inference—an operation that is inherently slow and occasionally unreliable due to stochastic failures (e.g., malformed outputs, null responses, or execution errors) (Ahn and Yin, 2025; Tan et al., 2025; Agrawal et al., 2024). To ensure scalability and robustness when processing large AIS data, we design a workflow management layer that automates task scheduling, anomaly handling, and de-redundancy. This layer includes two complementary components: the SD-KG Construction Workflow Manager and the Trajectory Imputation Workflow Manager.
4.4.1. SD-KG Construction Workflow Manager
The SD-KG Construction Workflow Manager coordinates a two–stage parallel workflow, comprising extraction and de-redundancy. It consists: Stack-Based Scheduler, Anomaly Guard, and Deredundancy Processor.
Stack-Based Scheduler. All minimal segments are ordered by timestamp and pushed as job tuples onto a compute stack $\mathcal{S}_{c}$, while a second stack $\mathcal{S}_{d}$ buffers validated micro-batches before de-redundancy. At runtime, the scheduler pops up to $b$ jobs from $\mathcal{S}_{c}$ to form a micro-batch $\mathcal{B}_{t}$ and dispatches them to $|\mathcal{B}_{t}|$ worker threads. Each worker executes the extraction pipeline to produce a per-segment knowledge unit $u=(v_{s},v_{b},v_{f})$, incrementing the retry counter $c$ on failure and discarding the job when $c>\varrho_{c}$. Validated outputs are packed as a micro-batch $\widehat{\mathcal{B}}_{t}$ and pushed onto the second stack $\mathcal{S}_{d}$, which acts as a decoupling buffer between extraction and de-redundancy. This design preserves temporal order and metadata while isolating the downstream consolidation stage from upstream variability.
Anomaly Guard. The generated knowledge unit $u$ may contain format errors or invalid content due to response timeouts, hallucinated fields, schema drift, or incomplete outputs. To protect the SD-KG, each produced unit is validated before any graph update. The validator checks that (i) the output is non-empty, (ii) the schema and format are correct, and (iii) all $(v_{s},v_{b},v_{f})$ components are present. Failed items are retried with bounded attempts: if $c<\varrho_{c}$, the job is re-queued with $c\!\leftarrow\!c{+}1$; otherwise, it is quarantined and excluded from subsequent stages. The validated subset of each batch is forwarded to the de-redundancy stack $\mathcal{S}_{d}$.
De-redundancy Processor. Before inserting new knowledge units into the SD-KG, redundancy elimination is applied to maintain compactness and prevent uncontrolled vocabulary growth. Validated batches are processed in parallel through: 1) behavior-token canonicalization merges semantically redundant tokens within behavior patterns using contextual and linguistic similarity analysis, ensuring that only canonical and representative tokens are retained across $\mathcal{P}^{s}$, $\mathcal{P}^{\theta}$, $\mathcal{P}^{\psi}$, and $\mathcal{P}^{i}$; 2) function-equivalence testing evaluates whether two imputation functions realize the same operational logic despite syntactic or parametric differences.
Finally, validated and de-duplicated knowledge units are written to the SD-KG. New nodes ($v_{s},v_{b},v_{f}$) are inserted if absent, and edge weights $(v_{s},v_{b})$, $(v_{b},v_{f})$ are incremented atomically to maintain statistical coherence.
4.4.2. Trajectory Imputation Workflow Manager
The Trajectory Imputation Workflow Manager mirrors the SD-KG Construction Workflow Manager but targets the reconstruction of incomplete minimal segments rather than graph construction. It omits the De-redundancy Processor, as imputation does not produce new knowledge units requiring de-duplication.
Stack-Based Scheduler. All incomplete minimal segments are timestamp-sorted and pushed onto a stack $\mathcal{S}_{i}$ in descending order. The scheduler repeatedly pops a batch of size $b$ and executes the imputation pipeline in parallel. Each job runs the three modules in Section 4.3 to reconstruct the missing segment using the SD-KG. A failure counter is maintained per job; when it exceeds the retry threshold $\varrho_{i}$, the job is aborted and logged for offline diagnosis.
Anomaly Guard. Imputation anomalies primarily arise from response timeouts or invalid model outputs. Each job therefore undergoes two basic checks: 1) the response is non-empty (i.e., no request timeout), and 2) the selected behavior pattern $v_{b}^{*}$ and imputation function $v_{f}^{*}$ are correctly retrieved from the SD-KG and are executable. Jobs failing either check are re-queued for retry until the per-item limit $\varrho_{r}$ is reached, after which they are quarantined for offline inspection.
Finally, the selected imputation methods are executed to reconstruct the missing trajectory segments. The resulting trajectories, along with their explanations, are recorded for interpretability and traceability. Further details (e.g., prompt about redundancy and overall procedures) are provided in Appendix A.3
Table 4. Performance of VISTA variants with different LLM configurations on AIS-DK and AIS-US datasets.
Analysis Programming Decision AIS-DK AIS-US
MAE(Lat) RMSE(Lat) MAE(Lon) RMSE(Lon) MHD MAE(Lat) RMSE(Lat) MAE(Lon) RMSE(Lon) MHD
Qwen-plus Qwen-plus Qwen-plus 1.817e-3 4.324e-2 1.123e-3 4.027e-2 0.2418 4.388e-3 1.455e-2 5.691e-3 2.052e-2 0.7945
Qwen-plus Qwen-plus Qwen-flash 5.706e-3 8.177e-2 5.578e-1 2.055e+1 2.8990 6.425e-3 2.957e-2 7.015e-1 1.311e+1 25.0557
Qwen-plus Qwen-flash Qwen-plus 2.872e-3 3.248e-2 5.818e-3 6.126e-2 0.5255 6.145e-3 2.371e-2 1.203e-2 6.505e-2 6.3462
Qwen-flash Qwen-plus Qwen-plus 1.902e-3 5.826e-2 1.216e-3 4.718e-2 0.2672 5.422e-4 1.7793e-2 8.960e-3 3.598e-2 1.8324
Table 5. Ablation study of the workflow management layer.
Batch Size AISDK AISUS
Vanilla w/o DR Vanilla w/o DR
Time Size Time Size Time Size Time Size
8 13:10:21 1,016 10:50:22 1,248 12:06:39 1,882 10:07:29 3,406
16 6:32:37 1,105 5:27:39 1,400 6:01:11 2,057 5:10:26 4,109
5. Experimental Study
5.1. Settings
5.1.1. Datesets
We use two AIS datasets: AIS-DK from the Danish Maritime Authority (Danish Maritime Authority, 2024) and AIS-US from the National Oceanic and Atmospheric Administration (National Oceanic and Atmospheric Administration, 2024). AIS-DK covers Danish waters in March 2024, including major routes in the Baltic and North Seas, with 10,000 vessel sequences and 2,000,000 AIS records from 348 vessels, each lasting on average 0.5 hours. AIS-US covers U.S. coastal waters in April 2024, with dense traffic near major ports, also containing 10,000 vessel sequences and 2,000,000 records from 4,723 vessels, each lasting on average 2.8 hours. Both datasets include various vessel types (e.g., Cargo, Tanker, Passenger), providing comprehensive coverage for evaluation.
5.1.2. Evaluation Metrics
For longitude and latitude, we report axis-wise Mean Absolute Error (MAE), a robust measure of average error, and Root Mean Squared Error (RMSE), which emphasizes large deviations, both expressed in degrees. For the joint spatial error, we use the Haversine distance (Wikipedia contributors, 2025b) with Earth’s mean radius R=6371R{=}6371 km and report the Mean Haversine Distance (MHD) in kilometers.
5.1.3. Baseline Methods
We compare VISTA with rule-based, deep learning-based, and LLM-based baselines. For rule based methods we include: Lin-ITP (Widyantara et al., 2023) linearly interpolates between anchors, Akima spline (Zaman et al., 2023) yields visually smooth, $C^{1}$-continuous paths that capture gentle curvature, and Kalman Filter (Wang et al., 2023) assumes a linear Gaussian state space with constant velocity and suits near linear motion. For deep based methods we include: MH-GIN (Liu et al., 2025) models multi scale features and cross attribute dependencies on a heterogeneous graph to impute all attributes in AIS data, and Multi-task AIS (Nguyen et al., 2018) is a recurrent framework with latent variables and AIS embeddings for noisy, irregular sampling. For LLM based methods we include: KAMEL (Musleh and Mokbel, 2023) treats completion as masked infilling with spatially aware tokenization; we also assess out of the box general models by using two families, Qwen (Qwen Team, 2025) and GLM 4.5 (Zeng et al., 2025), each with a lightweight and a full capacity variant with thinking mode enabled (Alibaba Cloud Team, 2025), yielding two pairs: Qwen-plus-th vs Qwen-flash-th and GLM-4.5-air-th vs GLM-4.5-th.
5.1.4. Experimental Settings
All experiments are conducted on a server with Intel Xeon Processor (Icelake) CPUs, 100GB RAM, and two NVIDIA A10 GPUs (each with 23GB memory). For data processing, each dataset is partitioned into 80% training and 20% testing. For comparability, all results reported below are computed on the held-out test split. To emulate realistic block missingness, trajectories are divided into minimal segments of length $m=20$, and each minimal segment is removed with probability $0.2$ (see Section 3.3). The mask configuration are shared for all methods. The more detailed experimental setting is provided in Appendix B.
Figure 7. Case study of VISTA.
5.2. Performance Evaluation
5.2.1. Effectiveness Analysis.
As shown in Table 2, which reports the results on AIS-DK and AIS-US across MAE, RMSE, and MHD, boldface denotes the best performance, while underlining indicates the second-best. Building on these results, we summarize four observations. 1) VISTA consistently outperforms all baselines, improving over the strongest baseline, MH-GIN, by 5%–94%. These results demonstrate its effective mining of AIS domain knowledge and productive use of underlying knowledge. 2) All models perform better on AIS-DK, and the inter-model spread is smaller than on AIS-US, consistent with shorter trajectory durations and smaller inter-sample time intervals in AIS-DK, which produce shorter missing gaps and an easier imputation setting. 3) Within the LLM group, the trajectory-oriented KAMEL performs markedly better than out-of-the-box general thinking-mode LLMs (Qwen-plus-th, Qwen-flash-th, GLM-4.5-th, GLM-4.5-air-th), underscoring the value of task-specific spatiotemporal tokenization and a “physical grammar” for trajectory imputation. Nevertheless, KAMEL still trails VISTA by a wide margin, highlighting the advantage of explicit, executable knowledge. The general models also show that “thinking” alone is insufficient, since capability strongly affects accuracy: higher-capacity variants outperform their lighter counterparts (Qwen-plus-th and GLM-4.5-th outperform Qwen-flash-th and GLM-4.5-air-th), and in our setting the GLM family tends to outperform Qwen (GLM-4.5-th and GLM-4.5-air-th outperform Qwen-plus-th and Qwen-flash-th), although all remain well below VISTA. 4) Deep-learning methods (Multi-task AIS, MH-GIN) generally outperform other kinds of baselines across metrics, reflecting the advantage of learning from AIS data; among them, MH-GIN is the strongest non-VISTA competitors overall, yet still trail VISTA by a wide margin.
5.2.2. Efficiency Analysis.
As shown in Table 3, VISTA attains the lowest time cost, reducing time by 51%–93% relative to the thinking-mode LLM baselines. The gains stem from two complementary factors: first, the SD-KG compresses raw AIS into structured priors that filter noise and constrain candidate behavior patterns and imputation functions, enabling targeted retrieval and lightweight execution instead of open-ended LLM reasoning; second, the Workflow Manager Layer coordinates parallel knowledge extraction and imputation with scheduling, fault tolerance, and de-redundancy, stabilizing stochastic LLM latency and eliminating redundant information in the SD-KG.
5.3. Ablation Study
We conduct two ablation studies. i) LLM role ablation: We vary the LLM capacity across three roles, Analysis (Behavior Abstraction and Explanation Composer), Programming (Method Builder), and Decision (Behavior Estimator and Method Selector), by downgrading one role while keeping the others fixed. Table 4 reports the results on AIS-DK and AIS-US. ii) Workflow ablation: We assess the workflow management layer under different batch sizes and with or without the De-redundancy (DR) mechanism. Table 5 presents the execution time and the total number of behavior pattern and imputation method nodes in the SD-KG (denoted as Size).
From Table 4, the Decision role is the most capacity-sensitive: downgrading it from Qwen-plus to Qwen-flash sharply increases MAE, RMSE, and MHD, showing that accurate behavior estimation and method selection demand stronger reasoning and contextual inference. In contrast, Analysis and Programming degrade more moderately under lighter models, as their semantic abstraction and method synthesis processes can still leverage the structured priors embedded in the SD-KG, which partially compensate for reduced model capacity. From Table 5, latency scales almost linearly with batch size, confirming that the workflow layer achieves efficient and contention-free parallel scheduling. Disabling de-redundancy slightly shortens runtime but significantly reduces size (e.g., $1,400\rightarrow 1,105$ on AIS-DK and $4,109\rightarrow 2,057$ on AIS-US), indicating a trade-off between speed and compact SD-KG.
5.4. Case Study
As shown in Figure 7, this case study illustrates VISTA’s knowledge-driven imputation in constrained waterways near Delaware Bay. A tanker with missing AIS data (1050–1100 s) was analyzed using structured data-derived knowledge and LLM reasoning. Based on preceding and following motion cues, VISTA identified a stable-speed, gradual port-side turn pattern, consistent with 67% of similar tankers “under way using engine” in this region. The selected spatial imputation function reconstructs the missing path as a smooth curved arc, preserving kinematic continuity. The human-friendly explanation links the maneuver to navigation in constrained waterways in New Castle County, Delaware, within designated corridors (Delaware River and Bay) governed by traffic separation schemes that require predictable motion aligned with channel geometry, and to operational protocols for large tankers that maintain stable speed and controlled heading, executing gradual turns (from south-southwesterly to southerly) for safe passage. The resulting trajectory aligns with maritime norms and environmental constraints, showing that the vessel’s behavior is compliant rather than anomalous, demonstrating how VISTA integrates data-driven evidence and domain reasoning for interpretable trajectory reconstruction.
6. Conclusion
We present VISTA, a knowledge driven and interpretable vessel trajectory imputation framework that returns reconstructed trajectory segments together with supporting knowledge for downstream use. VISTA defines this support as underlying knowledge that combines Structured Data-derived Knowledge (SDK) from AIS data with the implicit knowledge encoded in large language models (LLMs). VISTA organizes the SDK as a Structured Data-derived Knowledge Graph (SD-KG) and applies it in a data-knowledge-data loop. When going from data to knowledge, VISTA distills vessel attributes, behavior patterns, and validated imputation functions from raw AIS data and inserts them into the SD-KG. When going from knowledge to data, it uses this structured knowledge to estimate behavior, select and execute imputation methods, and generate human readable explanations grounded in maritime rules and operations. In addition, to ensure efficient processing of large scale AIS data, a workflow manager layer enables parallel scheduling, anomaly guards, and de-duplication, which reduces LLM latency and stabilizes throughput. Experiments on two AIS datasets show state-of-the-art accuracy and substantial efficiency gains, and the produced knowledge cues enhance downstream analysis; ablations show that the workflow management layer effectively reduces runtime through parallel scheduling and preserves the compactness of the SD-KG through the de-redundancy strategy.
In future work we plan to extend VISTA into a unified maritime trajectory management framework that supports multiple data management tasks, including trajectory imputation, anomaly detection, and trajectory prediction.
7. Acknowledgment
The research leading to the results presented in this paper has received funding from the European Union’s funded Project MobiSpaces under grant agreement no 101070279.
Appendix A The details of VISTA
A.1. The SD-KG Construction and Maintenance
Behavior Abstraction [TASK]
You are an expert in maritime data analysis.
Your task is to generate a list of specific, interpretable patterns that describe how both latitude and longitude (vessel positions) can be inferred from a set of AIS features.
These patterns will be used to impute missing values of latitude and longitude in AIS data. Each pattern must be:
Concrete and usable: describing a clear condition on the target features and the corresponding position range; Simultaneous: including both latitude and longitude ranges, or describing a trajectory pattern (e.g., a curved path, straight line, or repeated loop); Explainable, with a short justification after each pattern explaining why this condition relates to the given position; Mathematically expressive: including a possible trajectory equation or shape that approximates the vessel’s movement under this condition. [INPUT]
You are given by:
A sample of trajectory data (latitude, longitude, and various AIS features): {trajectory_data} [OUTPUT]
Please strictly follow the following format, return only one pattern,and output all patterns in triple quotes:
Pattern:
speed_pattern: speed profile without numerical values and punctuation (detailed description for speed profile). course_pattern: change in course over ground without numerical values and punctuation (detailed description for change in course over ground). heading_pattern: heading fluctuation without numerical values and punctuation (detailed description for heading fluctuation). intent: inferred maneuver intention without numerical values and punctuation (detailed description for inferred maneuver intention). ’’’ For speed_pattern, You can choose from {speed_dict}, and if you don’t have a suitable one, you can create a new one.
For course_pattern, You can choose from {course_dict}, and if you don’t have a suitable one, you can create a new one.
For heading_pattern, You can choose from {heading_dict}, and if you don’t have a suitable one, you can create a new one.
For intent, You can choose from {intent_dict}, and if you don’t have a suitable one, you can create a new one.
[EXAMPLE]
Pattern:
- speed_pattern: stable (the vessel is maintaining a consistent speed, not accelerating or decelerating)
- course_pattern: stable (the vessel is maintaining a consistent course over ground)
- heading_pattern: stable (the heading does not fluctuate significantly, indicating no sharp maneuvers)
- intent: navigating (the vessel is maintaining its course)
Make sure that your output strictly follows this format. Any patterns that do not adhere to this structure should be adjusted to fit the template. If any pattern involves multiple segments, please aggregate them.
Figure 8. Prompt of behavior abstraction.
A.1.1. Behavior Patterns Extraction
Figure 8 presents Behavior Abstraction Prompt, which operationalizes the behavior abstraction in Section 4.2.2 by mapping raw motion signals into a discrete, human-interpretable pattern for downstream imputation. The kinematic tokens $(p^{s},p^{\theta},p^{\psi})$ are instantiated as speed_pattern, course_pattern, and heading_pattern, respectively. The prompt explicitly instructs the LLM to analyze the temporal evolution within the provided {trajectory_data} and to choose from finite vocabularies ({speed_dict}, {course_dict}, {heading_dict}) before creating any new entries, thereby constraining vocabulary growth and enabling deduplication of semantically equivalent labels. And the intent token $p^{i}$ is realized through two complementary fields: a high-level intent (e.g., ‘navigating’) and a context-grounded navigation intent (e.g., ‘the vessel is maintaining its course’). Both are inferred from static and contextual cues, navigation status $\eta$, vessel type $\kappa$, and cargo type $\chi$ and spatial context $\sigma$ (obtained from Static and Spatial Encoder) in trajectory_data; a finite set {intent_dict} regulates label usage to preserve interpretability and retrieval efficiency. Finally, the output schema mandates a single structured pattern enclosed in triple quotes, ensuring robust parsing and consistent downstream indexing.
Imputation Method Builder [TASK] You are an expert in maritime trajectory analysis and spatial-temporal modeling, specialized in developing interpretable algorithms for vessel movement reconstruction. You are given vessel trajectory data and are asked to generate a spatial_function that estimates missing latitude and longitude positions based on known trajectory features and motion patterns. You need to adhere requirements: The spatial_function should be a single-line, directly executable Python function. The spatial_function is encouraged to choose from a variety of path models, not just linear interpolation. The spatial_function must compute a sequence of intermediate points using the provided start, end and Time_interval. start: A tuple representing the geographic coordinates (latitude, longitude) immediately before the missing data block.
Type: Tuple[float, float]
end: A tuple representing the geographic coordinates (latitude, longitude) immediately after the missing data block.
Type: Tuple[float, float]
Time_interval: A list of time differences in seconds relative to the timestamp of the starting point, covering the entire range including the point before and after the missing block.
Type: List[float]
[INPUT] You are given by: trajectory: {trajectory_data}. behavior pattern of trajectory: {pattern}. [OUTPUT] Please strictly follow the following format:
Function: ’’’ def spatial_function(start, end, Time_interval): return [...] ’’’
Description: A brief explanation of what this function does, including how it uses the input parameters. [EXAMPLE]
’’’def spatial_function(start, end, Time_interval): return []’’’
{feedback_text_description}
Figure 9. Prompt of imputation method builder.
[…]
’’’ Selected Movement ID: <ID>
Graph Support: <edges and weights you rely on>
Contextual Justification: <why consistent with boundary context>
Figure 10. The prompt of behavior selection and rationale.
A.1.2. Imputation Method Builder
Figure 9 illustrates the prompt design of the Imputation Method Builder for imputation method generation in VISTA. It instantiates the imputation method $v_{f}=(f,d(f))$ defined in Section 4.2.3 by compelling the LLM to output a single-line, directly executable Python function $f$ together with a concise textual description $d(f)$. The prompt enforces a fixed, machine-readable function format and explicitly defines the meaning of each input variable, where start and end denote the boundary coordinates of the missing segment and Time_interval specifies the temporal span across both sides of the gap. Contextual grounding is provided through trajectory_data and the behavior pattern pattern, guiding the generator to incorporate pattern-specific motion characteristics during function synthesis. By explicitly allowing non-linear paths (“not just linear interpolation”), the prompt expands the hypothesis space while maintaining direct executability and reproducibility via a unified output schema (triple-quoted code block plus a separate description).
Input: AIS dataset $\mathcal{X}$; thresholds $m$, $\varrho_{f}$ and $\varrho_{r}$.
Output: $\mathcal{G}_{d}=(\mathcal{V}_{s},\mathcal{V}_{b},\mathcal{V}_{f},\mathcal{E}_{sb},\mathcal{E}_{bf})$, $\mathcal{U}_{d}$
1 Initialize empty $\mathcal{G}_{d}$ and $\mathcal{U}_{d}$
2 foreach each vessel $\iota$ in $\mathcal{X}$ do
3  $\langle\mathbf{S}_{\iota}^{1},\dots,\mathbf{S}_{\iota}^{K}\rangle\;\leftarrow\;\textsc{Partition}(\mathbf{X}_{\iota};m)$
4  $\mathbf{M}_{\iota}\;\leftarrow\;\textsc{GetSegmentMask}(\langle\mathbf{S}_{\iota}^{1},\dots,\mathbf{S}_{\iota}^{K}\rangle)$
5  $\mathcal{U}_{d,\iota}\leftarrow\langle\rangle$
6  foreach $k=1$ to $K$ do
7     if $M_{\iota}^{k}=0$ then
8        $\mathcal{U}_{d,\iota}\leftarrow\mathcal{U}_{d,\iota}\|\varnothing$
9        Continue
10
11     /* Extract Static Attributes */
12     $\iota,\eta,\chi,\kappa\;\leftarrow\;\textsc{Mode}(\{x.\iota,x.\eta,x.\chi,x.\kappa\;|\;x\in\mathbf{S}_{\iota}^{k}\})$
13     $\tilde{d},\tilde{\ell},\tilde{\beta}\;\leftarrow\;\textsc{Discrete}(\{x.d,\ x.\ell,\ x.\beta\;|\;x\in\mathbf{S}_{\iota}^{k}\ \})$
14     $\sigma\leftarrow\textsc{Mode}(\{\textsc{Overpass}(x.\lambda,x.\phi)\;|\;x\in\mathbf{S}_{\iota}^{k}\})$
15     $v_{s}\leftarrow\{\iota,\eta,\chi,\tilde{d},\tilde{\ell},\tilde{\beta},\sigma,\kappa\}$
16     /* Extract Behavior Pattern) */
17     $p^{s},p^{\theta},p^{\psi},p^{i},p^{\tau}\leftarrow\textsc{BehaviorAbstraction}(\{x\;|\;x\in\mathbf{S}_{\iota}^{k}\},\mathcal{P}^{s},\mathcal{P}^{\theta},\mathcal{P}^{\psi},\mathcal{P}^{i})$
18     $v_{b}\leftarrow(p^{s},p^{\theta},p^{\psi},p^{i},p^{\tau})$
19     /* Generate Imputation Method */
20     $v_{f}\leftarrow\textsc{RetrieveBest}(v_{s},v_{b},\mathcal{V}_{f})$
21     if $v_{f}=\varnothing$ then $v_{f}\leftarrow\textsc{GenFunc}(v_{s},v_{b})$ ;
22     foreach $r=1$ to $\varrho_{r}$ do
23        $e\leftarrow\textsc{EvalMAE}(v_{f}.f,\mathbf{S}_{\iota}^{k})$
24        if $e\leq\varrho_{f}$ then break ;
25        else $v_{f}\leftarrow\textsc{GenFunc}(v_{s},v_{b})$ ;
26
27    /* Update SD-KG */
28     $\mathcal{V}_{s},\mathcal{V}_{b},\mathcal{V}_{f}\leftarrow\mathcal{V}_{s}\cup v_{s},\mathcal{V}_{b}\cup\{v_{b}\},\mathcal{V}_{f}\cup\{v_{f}\}$
29     $\mathcal{E}_{sb},\mathcal{E}_{bf}\leftarrow\textsc{EdgeInc}((v_{s},v_{b},v_{f}),\mathcal{E}_{sb},\mathcal{E}_{bf})$
30     $\mathcal{U}_{d,\iota}\leftarrow\mathcal{U}_{d,\iota}\|(v_{s},v_{b},v_{f})$
31
32 $\mathcal{U}_{d}\leftarrow\mathcal{U}_{d}\bigcup\{\mathcal{U}_{d,\iota}\}$
33
34return $\mathcal{G}_{d}=(\mathcal{V}_{s},\mathcal{V}_{b},\mathcal{V}_{f},\mathcal{E}_{sb},\mathcal{E}_{bf})$ , $\mathcal{U}_{d}$;
Algorithm 1 SD-KG Construction and Incremental Update
A.1.3. Overall Procedure of SD-KG Construction and Maintenance
Algorithm 1 summarizes the end-to-end pipeline that builds and updates the SD-KG from complete minimal segments. Line 1 initializes the empty graph $\mathcal{G}_{d}$ and the global per-vessel unit collection $\mathcal{U}_{d}$. Lines 2–5 segment each vessel’s stream (Partition) and build the mask (GetSegmentMask); then initialize the per-vessel knowledge unit sequence $\mathcal{U}_{d,\iota}\leftarrow\langle\rangle$, which records, in segment order, the knowledge unit built for each segment. Lines 6–9 skip incomplete segments: if $M_{\iota}^{k}=0$, append $\varnothing$ to $\mathcal{U}_{d,\iota}$ as a placeholder for a missing unit and continue; this preserves alignment between segment index $k$ and entries in $\mathcal{U}_{d,\iota}$. Lines 10–15 extract static attributes: Mode selects segment-wise majorities for categorical fields, Discrete bins draught and geometry (Section 4.2), and Overpass yields the spatial context; these form $v_{s}$. Lines 16–17 obtain the behavior pattern (Appendix A.1.1); together they form $v_{b}$. Lines 18–24 pick an imputation method $v_{f}$: RetrieveBest reuses a candidate from $\mathcal{V}_{f}$ using context and edge statistics (Section 4.3); otherwise GenFunc (Appendix A.1.2) generates executable code $f$. The loop validates with EvalMAE (mean of latitude/longitude MAE) against $\varrho_{f}$, refining via GenFunc up to $\varrho_{r}$ times. Lines 25–28 integrate results into SD-KG: nodes are inserted or reused, and EdgeInc creates or increments $(v_{s},v_{b})$ and $(v_{b},v_{f})$ edges to support later retrieval; append the knowledge unit $(v_{s},v_{b},v_{f})$ to $\mathcal{U}_{d,\iota}$ at position $k$. Line 29 aggregates each vessel’s knowledge unit sequence into the global collection: $\mathcal{U}_{d}\leftarrow\mathcal{U}_{d}\cup\{\mathcal{U}_{d,\iota}\}$. Thus, $\mathcal{U}_{d}$ is a vessel-indexed collection of sequences where each $\mathcal{U}_{d,\iota}$ preserves segment order and may contain $\varnothing$ at masked positions. Line 30 returns the updated SD-KG $\mathcal{G}_{d}$ and the knowledge unit collection $\mathcal{U}_{d}$.
A.2. Knowledge Driven Trajectory Imputation
A.2.1. Behavior Pattern Selection and Rationale.
Figure 10 presents the prompt used for LLM-driven Behavior Pattern Selection and Rationale described in Section 4.3.1. This prompt guides the LLM to identify the most behavior pattern for a missing trajectory segment and to articulate the reasoning behind this choice.
[…]
Method Selector and Rationale [TASK] You are an expert in maritime spatial-temporal modeling and trajectory reconstruction, responsible for evaluating and selecting the most suitable spatial function for accurate AIS trajectory imputation. Please select the most suitable spatial function for imputing missing latitude and longitude.
You need to adhere following requirements:
Direction: Which function has proven most reliable for similar kinematic patterns? Direction: Does the function’s underlying model (e.g., linear, curved) logically match the identified movement pattern (e.g., curved, straight)? Direction: Which function works best across different but related movement patterns? Direction: How well does each function handle the specific speed/course/heading characteristics? Important: When providing Statistical Support, ONLY discuss statistical evidence — DO NOT mention any edge-weights and graph but you could turn it to statistical describe. Important: Based on the calculation of weight proportions, all probabilities must be supported by evidence and cannot be arbitrarily fabricated. Each probability should be followed by its calculation process. Important: Avoid repeating, movement pattern analysis — focus on function execution quality. [INPUT] You are given by: Induced subgraph in DOT (Interpret weights as association frequencies): {dot_text} Functions with detailed information: {functions_text} behavior pattern that may correspond to missing parts: {movement_text} AIS data of the neighboring segments: {rows_text} [OUTPUT] Please strictly follow the following format:
Selected Function ID: <ID>
Statistical Support: <Don’t describe the edge weight, Don’t describe the graph support but you could turn it to statistical describe.1. Introduce the probability that this function can solve the missing value problem corresponding to the behavior pattern(Based on the calculation of weight proportions, all probabilities must be supported by evidence and cannot be arbitrarily fabricated. Each probability should be followed by its calculation process.). 2. Introduce the characteristics of this function and its degree of matching with the current context.>
Reasoning: <why this function technically fits the kinematic requirements>
Figure 11. Prompt of method selector and rationale.
Explanation Composer [TASK] You are an expert in maritime behavior interpretation and regulatory reasoning, specialized in translating computational decisions into human-understandable explanations for vessel trajectory analysis. Produce a human-friendly explanation for the chosen behavior and method. You need to adhere following requirements: Do NOT mention any node IDs or labels such as “Movement_Pattern_*” or “vessel_*”. Refer ONLY to the concrete attributes and descriptions provided below. [INPUT] You are given by: Induced subgraph in DOT: {dot_text} Selected movement (expanded): {movement_desc} Selected imputation method (expanded): {function_desc} Contextual vessel attributes (expanded list): {vessels_desc_block} Contextual behavior pattern: {vessels_behavior_pattern} [OUTPUT] Please strictly follow the following format:
Regulatory Rule Cue: <rule label + applicability + spatial anchor; leave Undetermined if insufficient>
Operational Protocol Rationale: <why this behavior is typical here; align with the spatial context; rule out key alternatives; do not mention IDs>
Figure 12. Prompt of explanation composer.
A.2.2. Imputation Method Selection and Rationale.
[…]
A strict output schema enforces machine-readability and interpretability: the model must return a triple-quoted block containing 1) Selected Function ID, 2) Statistical Support, which reports only statistical evidence (translated from association frequencies into probabilities) without mentioning graph structures or edge weights explicitly, each probability is accompanied by its calculation from weight proportions to prevent unverifiable claims—and and 3) Reasoning, a technical justification of why the method matches the kinematic requirements of $v_{b}^{*}$. This schema ensures the selected method $v_{f}^{*}\in\mathcal{C}_{f}^{K}$ is both verifiable (via explicit probability construction from $\mathcal{G}_{\iota,f}^{K}$) and actionable (via the prescribed imputation procedure), while the resulting rationale is stored in $\mathcal{J}_{\iota}^{k,f}$ for downstream auditing and integration.
De-redundancy Processor [TASK]
You are an expert in maritime knowledge consolidation and redundancy analysis, specializing in detecting and merging semantically equivalent vessel behavior patterns and imputation functions. You need to adhere following requirements:
For Behavior Pattern:
Exact duplicates: identical text Semantic equivalents: same meaning, different wording Minor variations: slight wording differences Overly specific terms: very detailed descriptions Contextual synonyms: same meaning in maritime context For Imputation Function: Functional equivalence: different implementations but same mathematical function Algorithmic similarity: same core algorithm with minor variations Parameter differences: same logic with different parameter values Code restructuring: same functionality with different code structure [INPUT]
You are given by:
Behavior patterns to analyze:
{vb_data_text}
Spatial functions to analyze:
{vf_data_text}
[OUTPUT]
Please strictly follow the following format:
For behavior patterns:
BEHAVIOR_REDUNDANCY:
[attribute_name]:
- <primary_term1> | [<redundant_term1>, <redundant_term2>]
- <primary_term2> | [<redundant_term3>, <redundant_term4>]
KEEP_UNIQUE: [<term1>, <term2>, <term3>]
For spatial functions:
FUNCTION_REDUNDANCY:
- <primary_function_id_or_code> | [<redundant_function_id_or_code1>, <redundant_function_id_or_code2>]
- <primary_function_id_or_code> | [<redundant_function_id_or_code3>]
KEEP_UNIQUE: [<function_id_or_code1>, <function_id_or_code2>]
Focus on maritime behavior semantics and functional equivalence. Preserve meaningful distinctions.
Figure 13. The prompt of de-redundancy processor.
A.2.3. Underlying Case for Human-friendly Explanation.
Figure 12 presents the prompt used by the Explanation Composer in Section 4.3.3. The prompt supplies a compact evidence view consisting of: i) the induced subgraph that binds the selected behavior and method with contextual entities, $\mathcal{G}_{d}\!\left[\mathcal{V}_{s}(\iota)\cup\{v_{b}^{*}\}\cup\{v_{f}^{*}\}\right]$ (serialized as DOT, {dot_text}), ii) the expanded description of the selected behavior $v_{b}^{*}$ ({movement_desc}), iii) the expanded description of the selected imputation method $v_{f}^{*}$ ({function_desc}), iv) the static vessel attributes $\mathcal{V}_{s}(\iota)$ ({vessels_desc_block}), and v) the boundary behavior context $(v_{b}^{-},\,v_{b}^{+})$ that anchors spatial and operational intent ({vessels_behavior_pattern}). These inputs jointly instantiate the spatial context and the vessel profile needed for explanation.
The prompt is constrained not to reference internal node identifiers and to ground its explanation only in the concrete attributes provided by the inputs. It must produce a fixed, machine-readable output with two fields: 1) Regulatory Rule Cue: a minimal yet sufficient statement of the governing rule in context (“rule label + applicability conditions + spatial anchor”), activated by $\sigma$ and $\mathcal{V}_{s}(\iota)$ and consistent with $v_{b}^{*}$; and 2) Operational Protocol Rationale: a procedural account of why the observed behavior typically occurs here, aligned with the rule cue and $\sigma$, and briefly ruling out salient alternatives. The returned triple-quoted block (see Figure 12) is recorded as $\mathcal{J}_{\iota}^{k,h}$ for downstream auditing and integration, ensuring that symbolic cues (rules and spatial anchors) and operational logic are preserved in a human-friendly, verifiable form without exposing graph-internal identifiers.
Input: AIS dataset $\mathcal{X}$; SD-KG $\mathcal{G}_{d}$; knowledge units of complete AIS data $\mathcal{U}_{d}$; thresholds $m$ and $K$.
Output: Set of knowledge-supported imputation outcomes $\mathcal{R}$.
1 $\mathcal{R}\leftarrow\varnothing$
2 foreach vessel $\iota$ in $\mathcal{X}$ do
3  $\langle\mathbf{S}_{\iota}^{1},\dots,\mathbf{S}_{\iota}^{K}\rangle\leftarrow\textsc{Partition}(\mathbf{X}_{\iota};m)$
4  $\mathbf{M}_{\iota}\leftarrow\textsc{GetSegmentMask}(\langle\mathbf{S}_{\iota}^{1},\dots,\mathbf{S}_{\iota}^{K}\rangle)$
5  $R_{\iota}\leftarrow\langle\rangle$
6  foreach $k=1$ to $K$ do
7     if $M_{\iota}^{k}=1$ then
8        $R_{\iota}\leftarrow R_{\iota}\|\varnothing$
9        continue
10     /* Context Extraction */
11     Obtain contextual knowledge units $u_{\iota}^{-},u_{\iota}^{+}$ from $\mathcal{U}_{d}$
12     Infer $\mathcal{V}_{s}^{k}(\iota)$ from neighbors $(u_{\iota}^{-},u_{\iota}^{+})$
13     Extract $(v_{b}^{-},v_{b}^{+})$ from $u_{\iota}^{-},u_{\iota}^{+}$
14     /* Behavior Estimator */
15     $\mathcal{C}_{b}\leftarrow\{\,v_{b}\in\mathcal{V}_{b}\mid\sum_{v\in\mathcal{V}_{s}^{k}(\iota)}w_{sb}(v,v_{b})>0\,\}$
16     Compute $\pi(v_{b})$ by Eq. 10;   $\mathcal{C}_{b}^{K}\leftarrow\textsc{TopK}(\mathcal{C}_{b},\pi,K_{b})$
17     $\mathcal{G}_{\iota,b}^{K}\leftarrow\mathcal{G}_{d}[\mathcal{V}_{s}^{k}(\iota)\cup\mathcal{C}_{b}^{K}]$;   $\texttt{dot}_{b}\leftarrow\textsc{SeriDOT}(\mathcal{G}_{\iota,b}^{K})$
18     $\langle v_{b}^{*},\mathcal{J}_{\iota}^{k,b}\rangle\leftarrow\textsc{Behavior\_Select\_Ration}(\texttt{dot}_{b},\ v_{b}^{-},v_{b}^{+})$
19    /* Method Selector */
20     $\mathcal{C}_{f}\leftarrow\{\,v_{f}\in\mathcal{V}_{f}\mid w_{bf}(v_{b}^{*},v_{f})>0\,\}$
21     Compute $\pi(v_{f})$;   $\mathcal{C}_{f}^{K}\leftarrow\textsc{TopK}(\mathcal{C}_{f},\pi,K_{f})$
22     $\mathcal{G}_{\iota,f}^{K}\leftarrow\mathcal{G}_{d}[\{v_{b}^{*}\}\cup\mathcal{C}_{f}^{K}]$;   $\texttt{dot}_{f}\leftarrow\textsc{SeriDOT}(\mathcal{G}_{\iota,f}^{K})$
23     $\langle v_{f}^{*},\mathcal{J}_{\iota}^{k,f}\rangle\leftarrow\textsc{Method\_Select\_Ration}(\texttt{dot}_{f},\ v_{b}^{*})$
24    /* Execution */
25     $\widehat{\mathbf{S}}_{\iota}^{k}\leftarrow\textsc{Execute}(v_{f}^{*}.f,\ \mathbf{S}_{\iota}^{k})$
26    /* Explanation Composer */
27     $\mathcal{G}_{\iota}^{h}\leftarrow\mathcal{G}_{d}[\mathcal{V}_{s}(\iota)\cup\{v_{b}^{*}\}\cup\{v_{f}^{*}\}]$
28     $\texttt{dot}^{h}\leftarrow\textsc{SeriDOT}(\mathcal{G}_{\iota}^{h})$
29     $\mathcal{J}_{\iota}^{k,h}\leftarrow\textsc{Exp\_Composer}(\texttt{dot}^{h},\ v_{b}^{*},\ v_{f}^{*},\ \mathcal{V}_{s}(\iota),\ v_{b}^{-},v_{b}^{+})$
30    /* Assemble Result */
31     $R_{\iota}\leftarrow R_{\iota}\|\big(\widehat{\mathbf{S}}_{\iota}^{k},\;((v_{b}^{*},\mathcal{J}_{\iota}^{k,b}),\ (v_{f}^{*},\mathcal{J}_{\iota}^{k,f}),\ \mathcal{J}_{\iota}^{k,h})\big)$
32
33 $\mathcal{R}\leftarrow\mathcal{R}\cup\{R_{\iota}\}$
return $\mathcal{R}$
Algorithm 2 Knowledge-Driven Trajectory Imputation
A.2.4. Overall Procedure of Knowledge Driven Trajectory Imputation
Algorithm 1 summarizes the dataset-level inference workflow that imputes all gaps using the SD-KG as a knowledge substrate; the concrete pseudocode is given in Algorithm 2. The procedure iterates over vessels and their segmented streams, retrieves boundary knowledge for each gap, selects a behavior and a method with graph-supported rationales, executes the method to reconstruct the missing segment, and composes a human-friendly explanation.
[…]
A.3. Workflow Management Layer
A.3.1. The de-reduplication strategy of De-redundancy Processor
[…]
The prompt enforces a machine-readable output that mirrors these partitions. For behaviors, the block BEHAVIOR_REDUNDANCY lists pairs “primary term \$|\$ [redundant terms]” grouped by attribute (e.g., maneuver, route, intent), with residuals reported in KEEP_UNIQUE. For functions, FUNCTION_REDUNDANCY lists “primary function id/code \$|\$ [redundant ids/codes}]”, with non-mergeable items collected under KEEP_UNIQUE. This schema yields two canonical sets $(\widehat{\mathcal{B}}_{b},\widehat{\mathcal{B}}_{f})$ and explicit many-to-one provenance mappings $(\pi_{b},\pi_{f})$, ensuring that SD-KG construction remains compact while preserving meaningful distinctions required for downstream reasoning and execution.
Input: AIS dataset $\chi$; batch size $b$; thresholds $\varrho_{c}$, $\varrho_{r}$, $\varrho_{f}$, $m$.
Output: SD-KG $\mathcal{G}_{d}$; knowledge units $\mathcal{U}_{d}$.
1
2/* Build Job Stack (Stack-Based Scheduler) */
3 $\mathcal{S}_{c},\mathcal{S}_{d}\leftarrow\textsc{EmptyStack}(),\textsc{EmptyStack}()$
4 foreach vessel $\iota$ do
5  $\langle\mathbf{S}_{\iota}^{1},\dots,\mathbf{S}_{\iota}^{K}\rangle\leftarrow\textsc{Partition}(\mathbf{X}_{\iota};m)$
6  foreach segment $\mathbf{S}_{\iota}^{k}$ sorted by timestamp do
[…]
8
9/* Parallel Knowledge Unit Extraction */
10 while $\mathcal{S}_{c}\neq\varnothing$ do
11  $\mathcal{B}\leftarrow\textsc{PopBatch}(\mathcal{S}_{c},\ b)$
12  /* Launch Parallel Extraction Jobs */
13  parallel for $(\iota,k,\mathbf{S}_{\iota}^{k},c)\in\mathcal{B}$
14     $u_{\iota}^{k}\leftarrow\textsc{ExtractKU}(\mathbf{S}_{\iota}^{k},\varrho_{r},\varrho_{f})$
15 /* Anomaly Guard */
16  $\mathcal{B}_{u}=\varnothing$
17  foreach $u_{\iota}^{k}\in\{\,u_{\iota}^{k}\mid(\iota,k)\in\mathcal{B}\,\}$ do
18     if $\textsc{Anomaly\_Detect}(u_{\iota}^{k})$ then
19        if $c<\varrho_{c}$ then
[…]
21
22     else
23        $\mathcal{B}_{u}=\mathcal{B}_{u}\cup\{u_{\iota}^{k}\}$
24
[…]
26
27/* Parallel De-redundancy Processor */
28 $\mathcal{U}_{d}\leftarrow\varnothing$
29 while $\mathcal{S}_{d}\neq\varnothing$ do
30  $\mathcal{B}\leftarrow\textsc{PopBatch}(\mathcal{S}_{d},\ b)$
31  /* Launch Parallel De-redundancy Jobs */
32  parallel for $\mathcal{B}_{u}\in\mathcal{B}$
33     $\mathcal{B}_{u}^{d}\leftarrow\textsc{Deredund\_Processor}(\mathcal{B}_{u})$
34 $\mathcal{U}_{d}\leftarrow\mathcal{U}_{d}\cup\{\,\mathcal{B}^{d}_{u}\mid\mathcal{B}_{u}\in\mathcal{B}\,\}$
35
36/* Write to SD-KG */
37 foreach $(v_{s},v_{b},v_{f})\in\mathcal{U}_{d}$ do
38  $\mathcal{V}_{s},\mathcal{V}_{b},\mathcal{V}_{f}\leftarrow\mathcal{V}_{s}\cup v_{s},\mathcal{V}_{b}\cup\{v_{b}\},\mathcal{V}_{f}\cup\{v_{f}\}$
39  $\mathcal{E}_{sb},\mathcal{E}_{bf}\leftarrow\textsc{EdgeInc}((v_{s},v_{b},v_{f}),\mathcal{E}_{sb},\mathcal{E}_{bf})$
40
return $\big(\mathcal{G}_{d},\ \mathcal{U}_{d}\big)$
Algorithm 3 SD-KG Construction Workflow Manager
A.3.2. Overall Procedure of SD-KG Construction Workflow Manager
Algorithm 3 orchestrates knowledge extraction with a two–stage parallel workflow (extraction $\rightarrow$ de-redundancy) connected by stack-based scheduling and batch barriers. Lines 1–6 initialize the workflow by constructing two stacks: the compute stack $\mathcal{S}_{c}$ and the de-duplication stack $\mathcal{S}_{d}$. All minimal segments $\mathbf{S}_{\iota}^{k}$ are pushed to $\mathcal{S}_{c}$ in timestamp order, each as a job tuple $(\iota,k,\mathbf{S}_{\iota}^{k},c)$ with an initial retry counter c=0c{=}0. Lines 7–12 launch batched parallel extraction. While $\mathcal{S}_{c}$ is non-empty, a batch of at most $b$ jobs is popped and executed concurrently (ParFor), where each job calls ExtractKU to produce a knowledge unit $u_{\iota}^{k}$ for segment $\mathbf{S}_{\iota}^{k}$. A synchronization barrier follows each batch to ensure consistent collection of results. Lines 13–21 perform anomaly detection and bounded retry. Each extracted unit is examined by Anomaly_Detect. If an anomaly is found and the retry counter satisfies $c<\varrho_{c}$, the job is re-queued to $\mathcal{S}_{c}$ with counter incremented to c+1c{+}1. Otherwise, the unit is admitted into the valid batch $\mathcal{B}_{u}$. Jobs that exceed $\varrho_{c}$ retries are quarantined. The validated batch $\mathcal{B}_{u}$ is then pushed onto $\mathcal{S}_{d}$ for de-redundancy. Lines 22–29 execute the parallel de-redundancy stage. Each batch popped from $\mathcal{S}_{d}$ is processed in parallel by Deredund_Processor, which merges semantically equivalent behavior tokens and function implementations to produce de-duplicated results $\mathcal{B}_{u}^{d}$. All cleaned units are accumulated into $\mathcal{U}_{d}$, enabling scalable consolidation independent of extraction. Lines 30–33 commit the consolidated results to the SD-KG. Each $(v_{s},v_{b},v_{f})\in\mathcal{U}_{d}$ is inserted or reused within $\mathcal{V}_{s},\mathcal{V}_{b},\mathcal{V}_{f}$, and the corresponding edges $(v_{s},v_{b})$ and $(v_{b},v_{f})$ are updated via EdgeInc. Each per-vessel sequence $\mathcal{U}_{d,\iota}$ is maintained in temporal order to align with segment index $k$. Line 34 returns the updated SD-KG $\mathcal{G}_{d}$ and the complete set of de-duplicated knowledge units $\mathcal{U}_{d}$.
Batch size $b$ constrains concurrent jobs to balance throughput and stability; the dual-stack scheduling pattern decouples extraction and consolidation, forming a pipeline that overlaps compute with integration. The retry threshold $\varrho_{c}$ safeguards against unstable segments, while the refinement bound $\varrho_{r}$ inside ExtractKU ensures quality control of generated imputation methods.
1
2
3
4
Input: AIS dataset $\mathcal{X}$; SD-KG $\mathcal{G}_{d}$; knowledge units of complete AIS data $\mathcal{U}_{d}$; batch size $b$; retry limit $\varrho_{i}$; thresholds $m$ and $K$.
Output: Reconstructed trajectories with explanations $\mathcal{R}$.
5
6/* Build Job Stack (Stack-Based Scheduler) */
7 $\mathcal{S}_{i}\leftarrow\textsc{EmptyStack}()$
8 foreach vessel $\iota$ do
9  $\langle\mathbf{S}_{\iota}^{1},\dots,\mathbf{S}_{\iota}^{K}\rangle\leftarrow\textsc{Partition}(\mathbf{X}_{\iota};m)$
10  foreach gap segment $\mathbf{S}_{\iota}^{k}$ sorted by timestamp do
[…]
12
[…]
14 while $\mathcal{S}_{i}\neq\varnothing$ do
15  $\mathcal{B}\leftarrow\textsc{PopBatch}(\mathcal{S}_{i},\ b)$
16  /* Launch Parallel Imputation Jobs */
17  parallel for $(\iota,k,\mathbf{S}_{\iota}^{k},c)\in\mathcal{B}$
18     $\mathcal{R}^{k}_{\iota}\leftarrow\textsc{Traj\_Imputation}(\mathcal{G}_{d},\mathcal{U}_{d},(\iota,k,\mathbf{S}_{\iota}^{k},c),K)$
19
20 /* Anomaly Guard */
21  foreach $\mathcal{R}^{k}_{\iota}\in\{\,\mathcal{R}^{k}_{\iota}\mid(\iota,k)\in\mathcal{B}\,\}$ do
22     if $\textsc{Anomaly\_Detect}(\mathcal{R}^{k}_{\iota})$ then
23        if $c<\varrho_{i}$ then
[…]
25
26     else
27        $\mathcal{R}=\mathcal{R}\cup\{\mathcal{R}^{k}_{\iota}\}$
28
29
return $\mathcal{R}$
Algorithm 4 Trajectory Imputation Workflow Manager
A.3.3. Overall Procedure of Trajectory Imputation Workflow Manager
[…]
Appendix B Experimental Configuration
B.1. Datasets
We use two AIS datasets: AIS-DK from the Danish Maritime Authority and AIS-US from the National Oceanic and Atmospheric Administration. AIS-DK covers Danish waters in March 2024, including major routes in the Baltic and North Seas; it contains 10,000 vessel sequences and 2,000,000 records, with an average sequence duration of 0.5 hours across 348 vessels. AIS-US covers U.S. coastal waters in April 2024 with dense traffic near major ports; it contains 10,000 vessel and 2,000,000 records, with an average sequence duration of 2.8 hours across 4,723 vessels. Both datasets include diverse vessel types (e.g., Cargo, Tanker, Passenger), providing representative coverage for evaluation. To construct these datasets, we first downloaded the raw AIS data from the official sources and performed data cleaning to remove abnormal or incomplete records. We then partitioned each vessel trajectory—March 2024 for AIS-DK and April 2024 for AIS-US—into fixed-length segments of 200 records, and uniformly sampled 10,000 segments to build the evaluation corpus.
B.2. Hyperparameters and Implementations
Lin-ITP, Akima spline, and Kalman Filter are implemented in Python using pandas and pykalman. MH-GIN (Liu et al., 2025), Multi-task AIS (Nguyen et al., 2018), and KAMEL (Musleh and Mokbel, 2023) follow the authors’ default configurations. GLM-4.5, GLM 4.5-air, Qwen-Plus (snapshot 0112), and Qwen-flash (snapshot 2025-07-28) are accessed through the Aliyun Bailian API (Aliyun Teams, 2025) with thinking mode enabled.
Vessel Trajectory Imputation [TASK]
You are a professional maritime data analyst. Please predict the missing trajectory segment based on the known vessel trajectory points.
You must strictly adhere to the following requirements:
Output strict JSON array format containing {missing_length} coordinate points Each coordinate point in [longitude, latitude] format, keep 6 decimal places Ensure smooth and continuous trajectory that follows vessel movement patterns Consider temporal continuity and generate reasonable intermediate trajectory Do not add any explanatory text, only output JSON array [INPUT] Previous trajectory point (last known point): {prev_str} Next trajectory point (first known point after gap): {next_str} [OUTPUT]
Predict the intermediate missing {missing_length} trajectory points between these two points.
[EXAMPLE]
[[121.123456, 31.234567], [121.124567, 31.235678], ...]
Figure 14. Prompt of general large language model for vessel trajectory imputation.
B.3. Evaluation Metrics
We evaluate only timestamps inside minimal segments $\mathbf{S}_{\iota}^{k}$ that are imputed. A segment-level mask $M_{\iota}^{k}\in\{0,1\}$ identifies real gaps ($M_{\iota}^{k}=0$) or fully observed segments ($M_{\iota}^{k}=1$) from which we create synthetic gaps (see Definition 3.8). Inside each selected segment, an internal mask $\widetilde{\mathbf{m}}_{\iota}^{k}\in\{0,1\}^{m}$ flags the timestamps that require evaluation. Let the global index set be
(11) $\mathcal{I}=\big\{(\iota,k,j)\,:\,\widetilde{m}_{\iota,j}^{k}=1\big\},\quad|\mathcal{I}|~\text{its size},$
and denote ground-truth and imputed coordinates at index $(\iota,k,j)$ by $(\lambda_{\iota,k,j},\phi_{\iota,k,j})$ and $(\hat{\lambda}_{\iota,k,j},\hat{\phi}_{\iota,k,j})$, respectively.
We adopt axis-wise Mean Absolute Error (MAE) to capture the average magnitude of errors robustly and Root Mean Squared Error (RMSE) to emphasize large deviations; both are reported in degrees. The formulas are:
(12) $\mathrm{MAE}_{\lambda}$ $=\frac{1}{|\mathcal{I}|}\sum_{(\iota,k,j)\in\mathcal{I}}\left|\hat{\lambda}_{\iota,k,j}-\lambda_{\iota,k,j}\right|,$
(13) $\mathrm{MAE}_{\phi}$ $=\frac{1}{|\mathcal{I}|}\sum_{(\iota,k,j)\in\mathcal{I}}\left|\hat{\phi}_{\iota,k,j}-\phi_{\iota,k,j}\right|,$
(14) $\mathrm{RMSE}_{\lambda}$ $=\left(\frac{1}{|\mathcal{I}|}\sum_{(\iota,k,j)\in\mathcal{I}}\left(\hat{\lambda}_{\iota,k,j}-\lambda_{\iota,k,j}\right)^{2}\right)^{1/2},$
(15) $\mathrm{RMSE}_{\phi}$ $=\left(\frac{1}{|\mathcal{I}|}\sum_{(\iota,k,j)\in\mathcal{I}}\left(\hat{\phi}_{\iota,k,j}-\phi_{\iota,k,j}\right)^{2}\right)^{1/2}$
We use the Haversine geodesic distance as the joint spatial error and report it in kilometer. The formulas are:
(16) $d_{\iota,k,j}=2R\,\arcsin\!$ $(\sqrt{\operatorname{hav}(\Delta\phi_{\iota,k,j}^{r})+\cos\phi^{r}_{\iota,k,j}\cos\hat{\phi}^{r}_{\iota,k,j}\,\operatorname{hav}(\Delta\lambda_{\iota,k,j}^{r})}),$
(17) $\mathrm{MHD}$ $=\frac{1}{|\mathcal{I}|}\sum_{(\iota,k,j)\in\mathcal{I}}d_{\iota,k,j},$
where $\operatorname{hav}(x)=\sin^{2}(x/2)$; $R$ is Earth’s mean radius (6371 km); $\Delta\phi_{\iota,k,j}^{r}$ and $\Delta\lambda_{\iota,k,j}^{r}$ are the latitude and longitude differences in radians; $\phi^{r}_{\iota,k,j}$ and $\hat{\phi}^{r}_{\iota,k,j}$ are the ground-truth and imputed latitudes.
B.4. Baseline Methods
We compare VISTA with three classes of imputation functions that align with Table 1: rule based trajectory imputation, deep learning based trajectory imputation, and LLM based trajectory imputation.
For rule based trajectory imputation, we include Lin-ITP (Widyantara et al., 2023), Akima Spline (Zaman et al., 2023), and Kalman Filter (Wang et al., 2023). 1) Lin-ITP linearly interpolates longitude and latitude between boundary observations, implicitly assuming constant velocity and heading, and is simple and fast. 2) Akima Spline constructs a piecewise cubic polynomial using local slopes estimated from neighboring points, avoiding oscillations and overshooting common in standard cubic splines while maintaining smooth curvature continuity. It performs better on irregular or nonuniform motion patterns. 3) Kalman Filter treats motion as a linear Gaussian state space with constant velocity dynamics; forward filtering with backward smoothing recovers latent states and positions and is most effective when kinematics are near linear under moderate Gaussian noise.
For deep learning based trajectory imputation, we include MH-GIN (Liu et al., 2025) and Multi-task AIS (Nguyen et al., 2018). MH-GIN is a multi scale heterogeneous graph imputation network for AIS streams: it extracts multi scale temporal features per attribute while preserving heterogeneous update rates, then constructs a multi scale heterogeneous graph to capture cross attribute dependencies and propagates information to fill missing values. Multi-task AIS is a recurrent framework with latent variable modeling and an embedding of AIS messages designed for high volume, noisy, and irregularly sampled data; it jointly supports trajectory reconstruction, anomaly detection, and vessel type identification.
For LLM based trajectory imputation, we include KAMEL (Musleh and Mokbel, 2023), Qwen-plus-thinking, Qwen-flash-thinking, GLM-4.5-thinking, and GLM-4.5-air-thinking. KAMEL casts trajectory completion as a missing word problem with spatially aware tokenization and multi point masked infilling. Beyond KAMEL, to probe the out of the box capability of general LLMs on trajectory imputation without any task specific design, we adopt two state of the art families, Qwen (Qwen Team, 2025) and GLM 4.5 (Zeng et al., 2025). For each family we include a lightweight and a full capacity variant, and we enable thinking mode (Alibaba Cloud Team, 2025) throughout. This yields two paired comparisons under a unified setup: Qwen-plus-thinking vs Qwen-flash-thinking for capacity scaling, and GLM-4.5-air-thinking vs GLM-4.5-thinking for the trade off between latency and accuracy.
B.5. Prompt of General Large Language Model for Vessel Trajectory Imputation.
Figure 14 presents the baseline prompt used to elicit intermediate points for a missing trajectory segment from a general-purpose LLM. The prompt conditions the model on two boundary observations, previous point ${p}^{-}=[\lambda^{-},\phi^{-}]$ ({prev_str}) and next (first post-gap) point ${p}^{+}=[\lambda^{+},\phi^{+}]$ ({next_str}), and a target count $m$ of missing samples ({missing_length}). The model must output a strict JSON array of $m$ coordinates $\big([\lambda_{1},\phi_{1}],\ldots,[\lambda_{L},\phi_{m}]\big)$, each in [longitude, latitude] with six decimal places, without accompanying text. The constraints emphasize (i) smooth and continuous spatial progression consistent with vessel motion heuristics and (ii) temporal continuity for plausible interpolation between ${p}^{-}$ and ${p}^{+}$. The example block in Figure 14 specifies the exact output schema to ensure deterministic parsing and fair, method-agnostic evaluation of baseline LLMs.